### v1

# Inferring fine-grained migration patterns across the United States

Abstract
Fine-grained migration data illuminate important demographic, environmental, and health phenomena. However, migration datasets within the United States remain lacking: publicly available Census data are neither spatially nor temporally granular, and proprietary data have higher resolution but demographic and other biases. To address these limitations, we develop a scalable iterative-proportional-fitting based method which reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data. We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 which captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data. These estimates are highly correlated with external ground-truth datasets, and improve accuracy and reduce bias relative to raw proprietary data. We publicly release MIGRATE estimates and provide a case study illustrating how they reveal granular patterns of migration in response to California wildfires.
1 Introduction
Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena. Migration patterns illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; effects of the COVID-19 pandemic [8]; and political polarization [9].
But migration data within the United States have serious limitations. The most fine-grained freely available datasets track migration at the county level. While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level. For example, research on flood-risk-induced migration uses models at the much more granular Census block level [10], arguing that this spatial granularity is necessary due to the highly localized nature of flood risk [11]. Research on neighborhoods which are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [12]. Movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 [13]. Publicly available, county-level migration datasets render these patterns entirely invisible.
Proprietary migration datasets offer greater granularity. Data aggregators like Infutor [14] combine many data sources — including voter files, property deeds, credit files, and phone books [15] — to attempt to infer address histories for individual-level movers. Such datasets have been widely used because they are extremely temporally and spatially fine-grained [16, 4, 17, 18, 19]. Previous work also suggests these data do contain valuable signal which correlates with external datasets, as validated by cross-referencing the data with hurricanes or public housing foreclosures [16], and with marginal population counts in the region of interest [15]. However, these proprietary datasets have three major disadvantages. First, they are not publicly available, limiting their utility to researchers. Second, they require extensive computational pipelines, and substantial computational resources, to clean and map to standardized geographical areas to facilitate subsequent analysis. Third, and most fundamentally, they combine multiple imperfect data sources using proprietary algorithms, and thus contain noise and biases. For example, Infutor and other consumer record datasets have been shown to overrepresent higher-income and majority-group populations in some settings [20].
To address these limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates which combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data. To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm which repeatedly scales a noisy matrix to reconcile it with constraints [21], to reconcile the raw Infutor migration matrix with the more reliable Census constraints. We accommodate uncertainty in address mapping and incorporate data on births, deaths, and immigration. The output of our method is a set of yearly inferred migration matrices at the Census block group (CBG) level from 2010-2019 covering the 50 states and the District of Columbia. Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows which are currently available on the 5-year level, and 18 million times more granular than the state-state flows which are currently available on the 1-year level.
We comprehensively validate MIGRATE by comparing it to external data sources. We show that our estimates correlate well with fine-grained Census and ACS data on population counts and migration flows. We show that MIGRATE improves accuracy relative to raw Infutor data: in particular, it reduces the deviation from ground truth ACS county-to-county migration estimates by 90%. We document systematic demographic biases in the Infutor data, which overcounts rural, older, white, and home-owning populations, and show that MIGRATE almost entirely eliminates these biases. To illustrate the benefits of granular migration data, we provide a case study using our estimates to study migration in response to California wildfires, showing that MIGRATE reveals dramatic increases in out-migration which county-level data obscures. To provide a foundation for more precise migration research in the social, environmental, and health sciences, we release MIGRATE for other researchers at migrate.tech.cornell.edu.
2 Creating MIGRATE
Here we provide a high-level summary of our method for inferring fine-grained migration matrices by harmonizing Census and Infutor data. In the Appendix, we provide full details of processing the Infutor dataset (Appendix A), processing Census data (Appendix B), and harmonizing the two data sources (Appendix C).
Preprocessing the raw Infutor dataset
We first preprocess the raw Infutor data into yearly migration matrices. The raw Infutor data consists of sequences of addresses for individual people. We use this information to compute the number of people moving between each pair of geographic areas. In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior. We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Appendix Table 2.
Individuals with active records (any time in 2010-2019) 374,217,253
Individuals with active records in a given year (average across 2010-2019) 231,270,602
US population in 2010 309,327,143
US population in 2019 328,329,953
Address records per active individual (mean) 2.67
Address records per active individual (median) 2
Table 1: Summary statistics for raw Infutor data and Census data during the period of interest (2010-2019). We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period. Comparing the average yearly number of active individuals in the Infutor data to the Census population shows that Infutor under-counts the population. (Table 2 provides the number of active individuals in each year from 2010-2019). Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
Preprocessing Infutor requires mapping addresses to geographic areas. When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder [22, 23]. We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results. The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code. We are able to map 99.21% of all addresses in Infutor to CBGs. We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in Appendix A.
Harmonizing the Infutor and Census datasets
Having produced the raw Infutor yearly migration matrices $E^{(t)}$, we next reconcile them with more reliable but less granular Census data by rescaling selected entries of $E^{(t)}$. For example, to reconcile $E^{(t)}$ with Census county populations at time $t$, we multiply the rows corresponding to each county by a scaling factor to match the Census population in that county. We rescale to different Census datasets in sequence, rather than simultaneously, because the datasets are not totally consistent with each other.
Figure 1: MIGRATE estimates. We estimate annual migration flows between all pairs of Census block groups (CBGs) from 2010-2019. (a) Average MIGRATE estimates of out-migration rates across the entire United States. (b-c) MIGRATE estimates of out-migration rates within New York City. MIGRATE estimates reveal granular spatial patterns invisible in publicly available county-to-county data (inset plot (b)). Out-migration rates for CBGs with fewer than 100 people are omitted.
Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations; then the 1-year counts of movers and non-movers by state; then the 1-year state-to-state flows; and finally the 1-year county populations. The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them: while CBG populations and state flows are estimated by the American Community Survey (ACS) and often contain large sampling errors, county populations are estimated by the Population Estimates Program (PEP) using more precise large-scale administrative datasets. We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (see Appendix D.2). To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [21, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges (Appendix C.4). The resulting migration matrices constitute our MIGRATE estimates. Appendix C provides additional methodological details on these scalings.
Naive implementations of our harmonization algorithm impose prohibitive computation times due to the size of the matrices involved. We rely on the sparsity of the problem to dramatically reduce both memory and time requirements, allowing our procedure to converge within 2 hours using 16G of memory and 8 cores. Our general approach — i.e., rescaling submatrices to match Census constraints — can straightforwardly be adapted to accommodate other sources of Census data.
Figure 1 depicts the MIGRATE estimates. Figure 1a plots the average out-migration rates for all CBGs in the United States in the 2010-19 period, highlighting the spatial granularity of the estimates. This granularity often reveals important patterns which are invisible at the county-county level, as we illustrate by zooming in on New York City and comparing the county-level rates obtained from ACS flows (Figure 1b) to the CBG-level out-migration rates inferred by MIGRATE (Figure 1c). In Appendix E, we include an additional analysis showing how the socioeconomic stratification of movers’ destinations changes depending on whether we look at county-level estimates or MIGRATE estimates: residents from lower-income CBGs in higher-income counties tend to move to other lower-income CBGs, but this trend is obscured at the county level.
3 Validating MIGRATE
We conduct extensive validations of the MIGRATE estimates. We first show that MIGRATE estimates are highly concordant with Census datasets and improve agreement relative to raw Infutor data (Figure 2). We then investigate demographic biases in the raw Infutor estimates and show that MIGRATE reduces them (Figure 3).
Figure 2: Validating the MIGRATE estimates. MIGRATE estimates are highly correlated with (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population). For in-migration rates, correlations are weighted by state or county population, and points are sized by population. (d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data. We compute RMSE between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates. Bars show the mean reduction in RMSE across all years; error bars plot standard deviation across years. For the in-migration rates, RMSE is weighted by area population.
MIGRATE estimates are highly correlated with ground-truth Census data
MIGRATE estimates are well-correlated with all available Census measures (Figure 2a-b). We first assess correlations with Census datasets incorporated in our harmonization procedure, to confirm it indeed helps reconcile MIGRATE with Census data. We would expect these correlations to be high but not uniformly perfect, because the Census datasets are not totally consistent with each other. By design, MIGRATE estimates perfectly match state and county populations (because the last step in our procedure is to match to county populations). MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations [^1] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data. We only use ACS data products whose period completely overlaps with the 2010-2019 MIGRATE range.. MIGRATE estimates are also highly correlated with state-to-state flows ($\rho=1$); this high correlation is driven in part by variation in state populations, as well as by the fact that most people do not move. Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix D.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
Importantly, we further verify that our estimation procedure yields highly correlated estimates even with datasets which are not used in our estimation procedure. This constitutes a validation on held-out data, a standard check for statistical estimation methods [25]. Specifically, we compare our MIGRATE estimates to ACS 5-year county-to-county flows, which are not incorporated in our estimation, and find they are highly correlated with our estimates ($\rho=1$). This high correlation is driven in part by the fact that most people do not move; hence, in Appendix D.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$). Another contributor to the high correlation in county-county flows is variation in county size, so we separately compute correlations with county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) which normalize for county size. Correlations remain high (population-weighted $\rho=0.715$).
As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation: for example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census block group populations). See Appendix D.2 for full results.
Overall, these validations demonstrate that MIGRATE estimates are highly correlated with external ground-truth Census data. This is not merely because of variation in state or county size (since it remains true when examining in-migration rates, which normalize for state or county size, as well as when looking at the tract and CBG level). Nor is it due merely to the fact that most people do not move (since it remains true when examining correlations for movers specifically). Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets which are not used in our estimation procedure.
MIGRATE estimates reduce error relative to raw Infutor data
We show that MIGRATE estimates also increase agreement with Census datasets compared to using the Infutor data. We note that Infutor systematically undercounts the population (Table 1); to compare to Infutor as generously as possible, and in particular to ensure that MIGRATE does not reduce error due to trivial rescalings, we multiply each raw Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population prior to conducting these comparisons. Figures 2c and 2d report the reduction in error that MIGRATE estimates achieve relative to Infutor data. We compute root mean squared error (RMSE) between (1) MIGRATE estimates and ground-truth Census data and (2) Infutor data and ground-truth Census data, and report the reduction in RMSE from using MIGRATE estimates. We show average reduction in RMSE across all data releases between 2010 and 2019; error bars represent standard deviation across these releases.
MIGRATE estimates eliminate error in state and county populations because they are constructed to match county-level estimates. Relative to raw Infutor data, MIGRATE estimates reduce error in Census tract populations by an average of $85.9\%$ and error in CBG populations by $83.6\%$. MIGRATE estimates reduce error in state-to-state flows by $89.5\%$ and error in state in-migration rates by $83.3\%$. We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix D.1.
We again confirm that we still see these reductions in error on held-out data which is not used in our estimation procedure. Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$. Neither of these datasets was used in producing MIGRATE estimates. We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix D.2). Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data.
MIGRATE estimates reduce demographic biases present in raw Infutor data
Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates. Infutor data displays biases which MIGRATE estimates greatly reduce. (a) Average errors in county populations in Infutor data relative to Census data; orange denotes counties where Infutor underrepresents the population, and purple denotes counties where Infutor overrepresents it. MIGRATE estimates remove all county-level errors by construction. (b) Spearman correlation between county demographics (x-axis) and error in Infutor estimates (y-axis). Infutor’s error is correlated with racial, socioeconomic, and other demographic characteristics. (c) Comparison of demographic bias in Infutor (purple) and MIGRATE (green). Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data. MIGRATE greatly reduces biases for all demographic subgroups.
The Infutor data displays geographic and demographic biases (Figure 3). Specifically, it overrepresents populations in the counties in the northeastern United States while underrepresenting populations in the southwest (Figure 3a). [^2] For this analysis, as above, we multiply each Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population to avoid deviations due to trivial rescalings. Errors plotted are averages of county population errors across all 10 years of data from 2010-2019. MIGRATE estimates remove all such county-level errors by construction.
Errors in the Infutor data also correlate with demographics (Figure 3b). Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups. These biases are consistent with biases found in previous research [20], as we discuss in detail in Appendix D.3, and could propagate into biases in downstream analyses of inequality and other topics.
MIGRATE estimates reduce biases in Infutor data. Figure 3c compares the demographic biases of the MIGRATE estimates to the biases of the Infutor data. To quantify bias, we compare (1) the ground-truth number of people within each group (from Census data) to (2) the number of people within each group estimated from Infutor or MIGRATE, which we compute as $\sum_{i}p_{i}\cdot n_{i}$, where $i$ indexes CBGs, $p_{i}$ is the proportion of people in a CBG within a given group (e.g., the proportion of Black residents) based on Census data, and $n_{i}$ is the number of people in a CBG in Infutor or MIGRATE. This quantifies how much Infutor and MIGRATE overcount or undercount a given demographic group, relative to ground-truth Census data. The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases. We present analyses for additional demographic groups, finding similar results, in Appendix D.4.
4 Case study: migration in response to California wildfires
Figure 4: Migration in response to California wildfires. (a) Out-migration following the Camp fire (2018; top) and Tubbs fire (2017; bottom). Red boundaries plot fire perimeters; black lines plot county boundaries; CBGs are colored by domestic out-migration rate in MIGRATE. Out-migration rates exceed 50% in many CBGs within the fire perimeters. (b) Out-migration rates in different groups of CBGs over time according to MIGRATE estimates. Out-migration rates in the year after the fire are higher in CBGs within the fire perimeter (red line) than in groups of CBGs outside the fire perimeter (other lines), including those neighboring the fire perimeter, those in affected counties, or those within California. (c) Out-migration rates in ACS 5-year county-to-county data remain relatively constant over time.
We demonstrate the value of the MIGRATE estimates by using them to study migration in response to wildfires in California. Natural disasters, including wildfires, are known drivers of human mobility [3, 26]. There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area. In many US states, including California, wildfire risk increased from 2010-2019 [27], and is expected to further increase due to climate change [28]. Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters [29].
Post-wildfire migration estimates can inform disaster response and long-term planning in multiple ways. First, migration data is crucial for policymaking in the aftermath of a wildfire. It helps guide the allocation of housing, public health, and other resources to support displaced residents while minimizing strain on housing markets in receiving areas [30, 26, 31, 32]. Data on out-migration rates from affected areas can also inform decisions about whether and how to rebuild in fire-prone regions [33, 34]. Second, high-resolution migration data is important for future wildfire planning. As populations relocate in response to shifting wildfire risks, migration patterns can refine policymakers’ estimates of future risk and guide the regulation of housing and insurance markets [35, 36, 37, 38, 39]. While some households may move to reduce their exposure, others may become “trapped” in high-risk areas due to financial, social, or logistical barriers [40, 41, 42]. Identifying these communities through high-resolution migration data can help ensure that government support reaches those who need it most [43, 34].
We use high-resolution fire perimeter data from the California Department of Forestry and Fire Protection [44]. Fire perimeters are typically much smaller than county boundaries, suggesting that analyzing fire impacts may require the granularity of the MIGRATE estimates as opposed to the relatively coarse county-to-county flows. We analyze the two most destructive fires in California from 2010-2019 (as quantified by the number of structures lost to the fire): the Tubbs Fire in October 2017, and the Camp Fire in November 2018 (Figure 4). The Tubbs Fire occurred in Napa County and Sonoma County and destroyed at least 5,636 residential or commercial structures [45]. The Camp Fire, about a year later, destroyed over 18,804 structures and damaged nearly 1,000 more in the Northern California county of Butte [46] [^3] The Camp Fire remains the most destructive wildfire in California as of early 2025; the Tubbs Fire has only been surpassed by the January 2025 Eaton and Palisades fires [47]..
MIGRATE estimates reveal dramatic levels of out-migration for CBGs which lie within the fire perimeters (Figure 4a) in the year following the fires, often exceeding 50%. [^4] The fact that the out-migration rate is not 100% is likely due to a number of factors, including the often-preferred option of rebuilding and returning [37, 38]. Looking at administrative address data from mail delivery after Camp Fire, for example, researchers found a much more significant uptick in short-term vacancy than in long-term vacancy [29]. In contrast, CBGs which do not lie within the fire perimeters experience much lower out-migration rates. We systematically quantify these differences in Figure 4b, which compares CBGs within the fire perimeter to three sets of less-affected CBGs outside the perimeter: other CBGs neighboring the perimeter; other CBGs within the affected counties; and others within California as a whole. The out-migration rate in the year following the Camp Fire for CBGs within the fire perimeter is 46%, at least 3.1$\times$ that of less-affected CBG groups; the out-migration rate in the year following the Tubbs Fire for CBGs within the perimeter is 37%, at least 2.8$\times$ that of less-affected CBG groups. Our estimates of out-migration following the Camp Fire are similar to those in prior work [26].
In contrast, publicly available county-level migration data obscure these dramatic out-migrations; Figure 4c shows that rates of out-migration in affected counties remain essentially flat. This happens because the county-level data is both too spatially and too temporally coarse: the county boundaries include many CBGs unaffected by the fire, and the estimates are aggregated over 5 years. Further, many people displaced by the Camp and Tubbs Fires moved to other CBGs within the affected counties (and thus would not be counted as movers in the county-level data): 77% of movers from CBGs within the Tubbs fire perimeter and 54% of movers from CBGs within the Butte Fire perimeter remained within the affected counties.
ACS 5-year CBG population estimates similarly obscure the dramatic levels of out-migration due to their lack of temporal granularity. In the year following the fires, MIGRATE estimates reveal population declines 260% larger in magnitude for the Tubbs fire and 40% larger in magnitude for the Camp fire than those visible at the 5-year ACS level. (For this analysis, we compare to the relative change in the ACS 5-year dataset released the year following the fire from the ACS 5-year dataset the year before: for example, for the 2018 Camp Fire, we compare the 2015-2019 ACS estimates to the 2014-2018 ACS estimates.) The ACS 5-year population estimates, unlike MIGRATE estimates, also provide no insight into the destinations of out-movers.
Collectively, these analyses showcase how MIGRATE estimates reveal substantial, policy-relevant responses to wildfires which traditional data sources conceal. Indeed, the MIGRATE estimates enable further analyses beyond those conducted here: for example, they can illuminate inequality in responses to wildfires by connecting out-migration patterns with CBG-level demographics.
5 Discussion
We produce MIGRATE, a dataset of spatially and temporally fine-grained migration matrices capturing annual flows between pairs of CBGs from 2010-2019. Our estimates are approximately 4,600 times more spatially granular than publicly available county-to-county 5-year migration data, and 18 million times more spatially granular than state-to-state 1-year migration data. We produce these estimates by developing an IPF-based algorithm to harmonize granular but biased proprietary migration data with more reliable but less granular Census datasets.
We show that MIGRATE estimates correlate highly with external Census ground-truth datasets and reduce error and demographic bias relative to the raw proprietary data. We discuss the measures taken to protect the privacy of all individuals in the dataset in Appendix F. Our estimates are available at migrate.tech.cornell.edu.
While we validate MIGRATE estimates comprehensively against external data sources, practitioners making use of the data should be mindful of two limitations. First, while we use multiple ground-truth Census data sources to reduce the biases we document in the raw Infutor data, uncorrected biases likely remain. For example, while we correct biases in CBG populations, we cannot correct sub-CBG-level population biases, or biases in CBG-CBG flows. Second, while we show that MIGRATE estimates are highly correlated with external Census datasets (including held-out datasets which are not included in producing our estimates), these validations have imperfections. As we further describe in Appendix B.2, Census datasets themselves have biases; can possess significant margins of error, particularly at fine spatial scales; and are not perfectly consistent with each other. Another challenge in validation is the lack of a ground-truth CBG-CBG migration matrix against which to validate; publicly available Census datasets do not afford this level of granularity, and non-Census data sources (e.g., voter files) have significant biases and measure substantively different populations than the one we seek to model [48, 49]. Overall, MIGRATE should be viewed as a new data source which like all data possesses limitations but nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
We hope these improvements will facilitate a wide range of further migration-related analyses. As our wildfire case study illustrates, MIGRATE reveals patterns which remain invisible in publicly available migration data. The fine spatial scale of the data also enables analyses of inequality in migration, by linking migration patterns to CBG-level demographics. Our work also highlights methodological directions for improving estimation of fine-grained migration patterns, including providing principled margins of error on estimates and incorporating additional data sets to further improve estimation and validation. More broadly, we hope MIGRATE will facilitate more precise study of many migration-related phenomena across the social, health, and environmental sciences.
Appendix A Processing Infutor Data
Infutor provides data in tabular form. Each row in the data provides data for one individual, which includes a list of known addresses, along with the date when the individual is first observed at the address (which we refer to the effective date below), and the first and last date that the individual is observed in the data (which we refer to as the listed start date and the listed end date, respectively). [^5] For a subset of individuals, Infutor also provides estimated demographic information (e.g., age, gender, and socioeconomic information). We do not incorporate this information in our analysis for two reasons: first, complete demographic data is only available for about half of all individuals; second, the reliability of estimated demographic information is concerning—as the data lacks documentation, we should not discard the possibility that demographic information has been imputed based on addresses—and contains known biases [20]. Dates are listed as month and year. For example, the data for one individual might consist of two addresses: “1 Main Street, Everytown, USA, 10000 (January 2010); 2 Cornelia Street, New York, New York, USA (December 2017)” followed by an initial date of January 2008 and an end date of December 2017. We describe the steps taken to (1) identify and clean Infutor data records within the scope of MIGRATE, (2) process the Infutor data into matrices documenting yearly flows between address pairs, and (3) map these address-level matrices to the CBG-level matrices $E^{(t)}$ mentioned in the main text. Table 2 summarizes the raw and processed data.
A.1 Cleaning address histories
We first create a sequence of monthly addresses for each individual in Infutor: i.e., the address at which they are living each month. This requires us to resolve any inconsistencies in the dates provided by Infutor and model uncertainty in address histories.
We define the interval of activity during which each individual is active (observed) in the Infutor data. For each individual, we define their reconciled start date as the minimum of their first effective date, and their listed start date. We similarly define their reconciled end date as the maximum of their last effective date, and their listed end date. (For example, if the effective date list for an individual was [January 2013, January 2016], their listed start date was January 2014, and their listed end date was January 2017, their reconciled start date would be January 2013, and their reconciled end date would be January 2017.) Finally, we define their interval of activity as [reconciled start date -- 1 year, reconciled end date $+$ 1 year]. We use 1-year padding to reflect the annual granularity of the final dataset we produce and avoid discarding data in each yearly estimation process. This padding is also consistent with our treatment of deaths and emigrants, who will be recorded as non-movers in MIGRATE (and thus keep residence in the address they held during their last year alive in the United States).
Having defined each individual’s interval of activity, we define the address at which the individual is located for each month in that interval. This requires us to resolve inconsistencies in the address dates provided by Infutor, which we do as follows. We discard any addresses lacking an effective date, unless they are the only address for the individual—in such cases, we consider the reconciled start date also to be the effective date for that address. We discard any postal box addresses if there is a non-postal box address for an individual with an effective date within a year, since we assume that non-postal-box addresses are more reliable indicators of where someone lives. We then forward fill the remaining addresses so that each address spans span the period between its own effective date and the effective date of the next address. If the reconciled start date and reconciled end date for the individual differ from all the address effective dates, we fill these gaps with the first and last available addresses chronologically. If the individual has multiple addresses with the same effective date, we uniformly split the individual between the addresses by saying there was an equal probability of residence in any of them. At the end of this process, each individual is associated with a monthly probability distribution over addresses for each month in their interval of activity.
A.2 Processing address histories into annual address-address migration matrices
When then aggregate the monthly address histories for each individual to the annual level in a way which is consistent with the ACS interview process. We will ultimately reconcile the Infutor data with ACS geographic mobility data and thus want these datasets to be consistently processed. Our goal is to simulate the sampling process of ACS, in which any individual can be asked, throughout the year, where they lived one year ago (the specific wording of the ACS interview question is, “Where did this person live 1 year ago?”). To do so, we loop over the twelve months of the year and compute how the individual would answer this survey question if they were asked in each month, yielding flows between a pairs of addresses (where the individual currently resides, and where they resided a year prior).
We allow monthly flows to be probabilistic. For each pair of months in which the individual was certainly seen at single addresses, they would notify the ACS of that move (or stay) with probability one. However, if in one of the months there was uncertainty due to conflicting effective dates, the corresponding flow will also be uncertain; the individual might notify ACS they moved from any of the addresses they resided in the previous year during the month, or to any of the addresses they reside in the current year. We multiply probabilities of residence in each different address to obtain the probability of a flow; if the monthly address distribution remains constant in both years we assume permanence (i.e. the individual did not move) and report that individual possibly stayed in each listed address with its own residence probability. Accounting for this uncertainty yields a list with 3-tuples of addresses and a probability (ADDR1, ADDR2, $p_{1\rightarrow 2}$) per month; weighting these tuples per month (according to the number of days in the month) yields an yearly distribution.
We then aggregate these yearly distributions over the entire Infutor data, we create a list of expected ACS responses for the full population. That is, if multiple people reported a flow between ADDR1 and ADDR2, we add all their individual probabilities and report a single combined flow between ADDR1 and ADDR2. This represents the expected number of individuals with that flow reported to ACS. Finally, we create an yearly set of matrices $A^{(t)}$ with dimension $N_{\text{ADDRESSES}}\times N_{\text{ADDRESSES}}$ (i.e. a square matrix on the number of unique Infutor addresses) where each entry $A^{(t)}_{ij}$ corresponds to the expected number of individuals moving between addresses $i$ and $j$ from year $t-1$ to year $t$. These matrices will be further aggregated based on the location of each address.
A.3 Processing address-address migration matrices into CBG-CBG migration matrices
Figure 5: Flowchart detailing the process of mapping addresses to CBGs. We start with all Infutor addresses and classify them into five address types. We remove addresses which lie within US territories but not US states (around .26% of the addresses). Incomplete addresses—those that do not have a precise street address line—as well as PO boxes and rural routes are mapped probabilistically to CBGs according to their ZIP code. The vast majority (90.4%) of addresses are clean, complete street addresses which are sent to the Census geocoder, which is able to map 81.26% of these addresses to a Census block group. If the Census geocoder fails to provide a match, we reattempt matching using ESRI’s ArcGIS geocoder; this achieves a lower match rate of 58.24%, in part because the sample it is applied to is more difficult to parse. In total, we are able to map 99.21% of all addresses in the original Infutor data: 83.33% to a precise latitude and longitude, and 15.89% to a ZIP code.
We transform the address-level migration matrices $A^{(t)}$ into CBG-level matrices $E^{(t)}$ by mapping each address either to a precise CBG where possible, or to a distribution over CBGs when only a ZIP code is available. We use the 2010 Census block group boundaries for all addresses to maintain geographic consistency. Figure 5 details this process.
Our address mapping pipeline is divided in two parts: mapping addresses via geocoding, and mapping addresses with a ZIP code-to-CBG crosswalk. We initially attempt to use one of two state-of-the art geocoders to map addresses to latitudes and longitudes: the publicly-available Census Bureau geocoder [23] and ESRI’s ArcGIS geocoder [22]. We first attempt to geocode an address via the Census geocoder; if this algorithm does not find a match, we submit the address to ArcGIS. If neither algorithm finds a match, we map the address to Census Tracts intersecting its ZIP code, weighting probabilities based on the share of residential ZIP code addresses represented by the population of each Census Tract. Addresses representing rural routes and postal boxes are directly mapped via ZIP code, as well as incomplete addresses. We use the HUD ZIP-to-tract crosswalk, with the date closest available to the last seen date of an address to account for the fact that ZIP code definitions may vary and are not nested within Census geographies [50]. We then distribute the probability to each Census block group within the tracts, again weighting by their relative population shares. Success rates for each of these steps are high and detailed in Figure 5.
We use the results of this mapping process to create an auxiliary matrix $\mathcal{G}$ of dimensions $N_{\text{ADDRESSES}}\times N_{\text{CBGs}}$. Each row of $\mathcal{G}$ defines a probability distribution over Census block groups: an address $i$ belongs to CBG $j$ with probability $\mathcal{G}_{ij}$. If the address has been precisely mapped (i.e., geocoded via the Census or ArcGIS geocoders), the corresponding row of $\mathcal{G}$ has a single entry equal to $1$, with other entries 00. We use the matrix $\mathcal{G}$ to compute the CBG-to-CBG matrix $E^{(t)}$ via the following equation:
$E^{(t)}=\mathcal{G}^{T}\cdot\left[A^{(t)}-\text{diag}\left(A^{(t)}\right)\right]\cdot\mathcal{G}+\text{diag}\left[\mathcal{G}^{T}\cdot\text{diag}\left(A^{(t)}\right)\right]$ (1)
where the $\text{diag}\left(\cdot\right)$ operator either converts a vector into a diagonal matrix or extracts the diagonal of a matrix as a vector. The first term in this equation maps movers to the appropriate entries of the migration matrix: for example, a precisely mapped move from CBG $i$ to CBG $j$ will increment $E^{(t)}_{ij}$ by 1, and a move from ZIP code $a$ to ZIP code $b$ will distribute mass (summing to 1) evenly across the sub-block whose rows lie within ZIP code $a$ and columns lie within ZIP code $b$. The second term in this equation maps stayers to the appropriate diagonal entries of the migration matrix: for example, a precisely mapped stayer in CBG $i$ will increment $E^{(t)}_{ii}$ by 1, and someone who stays within ZIP code $a$ will distribute mass (summing to 1) evenly across the diagonal entries corresponding to CBGs within that ZIP code.
Table 2 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline. The number of individuals with active records in the data represents the number of individuals with an interval of activity containing each year at the start of the data processing; the number of CBG moves corresponds to the sum of off-diagonal entries of each matrix $E^{(t+1)}$ and measures the “effective size” of each year’s data: how many movers we actually account for at the beginning of the estimation procedure.
Year (1) Active records (2) US Population (3) Effective Dates (4) Processed CBG Moves
2010 269,228,776 309,327,143 48,225,002 15,180,982
2011 260,064,921 311,583,481 28,196,562 11,793,806
2012 259,451,204 313,877,662 32,009,822 12,544,199
2013 252,933,953 316,059,947 27,386,564 13,829,056
2014 256,453,630 318,386,329 37,386,684 13,071,070
2015 241,295,957 320,738,994 30,777,204 14,022,249
2016 242,298,262 323,071,755 38,085,337 13,605,804
2017 193,479,001 325,122,128 34,599,168 12,684,184
2018 177,629,604 326,838,199 33,183,982 13,505,890
Table 2: Yearly Breakdown of Data Summaries. Population and addresses accounted by the Infutor and Census datasets during the analysis period (2010-2019). (1) Individuals have active records in a given year if the year falls within their interval of activity in the dataset. The number of active individuals decreases in recent years, as Infutor records tend to appear when a move is reported. Comparison of these values to (2) the Census population shows that Infutor under-counts the population at every given year. (3) The number of effective dates per year is naturally reduced as we process the data into (4) CBG-to-CBG moves: an individual with two addresses will have at most one move, some addresses do not constitute residences, some individuals may have a single address in the year thus not move, etc.
Appendix B Processing Census Data
The U.S. Census Bureau runs multiple programs with the aim of counting and estimating demographic and socioeconomic population data. Each of these programs may release data at different spatiotemporal granularity and, even if geographies or timespans agree, the data might lack internal consistency. We use a series of steps to clean, select, and process Census datasets. We describe our procedures for reconciling Census geographies (§B.1), selecting which Census datasets to use as constraints (§B.2), and how we process the Census data to deal consistently with births, deaths, emigration, and immigration (§B.3).
B.1 Reconciling Census geographies
The United States is hierarchically divided into states, counties, Census tracts, Census block groups, and Census blocks, with each level subdividing its parent geography. Geographic boundaries are not necessarily consistent over time, and ensuring that population counts reflect the same geographic area across multiple years is essential when working with longitudinal data. For example, to produce multi-year estimates, the American Community Survey maps every reported address to the corresponding geography in the current year [51]. We use the 2010 Census boundaries for all addresses to maintain geographic consistency.
While the vast majority of geographies remain intact in the inter-decennial period (between Censuses, 2010-2020), there are some geography changes we must address. When geographies merge or split throughout the decade, we resolve the change by keeping only the coarser geographic boundary — i.e., the combined area which is the union of all finer-grained areas. With this approach, we can aggregate population statistics from the fine-grained areas into the coarse areas. There was one such county change in 2010-2019 affecting Bedford County in Virginia [52] and a few Census tract or Census block group changes documented yearly in the Census Bureau website (e.g. [53] for 2010 to 2011). When we aggregate ACS estimates that contain margins of error, we also aggregate margins of error in the L2 norm—as proposed by the ACS [54]. For our purposes, area renames are irrelevant. After these adjustments, MIGRATE considers a universe of $217,740$ Census block groups grouped into $3,142$ counties.
B.2 Selecting Census datasets
2010 2011 2012 2013 2014 2015 2016 2017 2018 2019
CBG Populations 15.76 15.67 15.42 15.33 15.17 14.94 14.95 15.08 15.21 15.59
Tract Populations 6.41 6.24 6.03 5.95 5.79 5.60 5.50 5.54 5.56 5.66
County Flows - 91.03 89.97 89.47 89.41 88.06 87.33 87.30 87.01 87.01
County Non-Movers - 1.78 1.72 1.69 1.68 1.65 1.64 1.65 1.63 1.65
County In-Migration Rate - 16.37 15.79 15.55 15.48 15.03 14.96 15.25 15.34 15.60
State Flows - 47.61 45.47 45.14 44.23 45.33 45.76 46.31 45.62 47.31
State Non-Movers - 0.41 0.39 0.39 0.38 0.40 0.40 0.38 0.37 0.39
State In-Migration Rate - 4.95 4.68 4.79 4.65 4.70 4.76 4.72 4.85 5.02
Table 3: ACS Coefficients of Variation. Average coefficient of variation (in %) of the non-zero estimates in each dataset. CVs were computed as the ratio of the standard error (obtained by dividing the margin of error by $1.645$) by the estimate [54]. We did not use flows released in 2010 (as they would report moves from 2009).
We rely primarily on three Census datasets: the ACS 5-year CBG populations; ACS 1-year state-to-state flows (from which we use raw flows as well as counts of non-movers and movers); and PEP 1-year county populations. PEP makes use of administrative records on births, deaths, and net migration to directly estimate county-level populations; the numbers reported contain the population estimate for July 1st of every year. ACS estimates, on the other hand, are a product of year-round surveying aggregated at the correspondent time scale, then released with associated margins of errors that define $90\%$ confidence intervals [54]. Our dataset selection and methodology accounts for these different precision levels across the board.
We would like to avoid matching to datasets that were too noisily estimated. To decide which ACS datasets to use in our harmonization process, we examine their level of sampling error. We quantify the sampling error by the coefficient of variation (CV): the ratio of the standard deviation to the estimate value. Table 3 reports the mean coefficient of variation in ACS estimates. In general, the number of non-movers has very low CVs (mean below 2% for counties and below 1% for states); population estimates have considerably lower CVs than flow estimates; state-to-state flows have relatively high CVs (mean around 45% every year), and county-to-county flows have extremely high CVs (mean almost 90%). We opt not to match to county-county flows due to the imprecision of raw flow estimates, although we do use them as a validation. In-migration rates have much tighter CVs than the corresponding flows (mean around 15% and 5% for counties and states, respectively), and we also use those as validations.
We perform a synthetic simulation study to verify that the datasets we do match to — state-to-state flows, and CBG population estimates — have sufficiently low sampling error to be reliably used. In particular, we resample each dataset using its reported margins of error, and compute correlations between the resampled data and the original data. To also assess mobility rates, we summarize state-to-state flows via the net migration rate. CBG populations and net state migration rates remain highly correlated with themselves after resampling (average Pearson correlation of $\rho=0.976$ and $\rho=0.766$, respectively).
Based on these analyses, we harmonize the entries of each yearly matrix $E^{(t)}$ with the following data: (1) CBG populations (both from 2010 Census and 5-year ACS); (2) ACS 1-year counts of movers and non-movers by state; (3) ACS 1-year state-to-state flows; and (4) PEP 1-year county populations.
B.3 Accounting for natural population increase and international migration
Using Census datasets to constrain MIGRATE estimates requires ensuring that the Census datasets reflect the same population accounted for in the Infutor matrix $E^{(t)}$. Each entry $E^{(t)}_{ij}$ represents the expected number of people alive with a residence in area $i$ at year $t-1$ and with a residence in area $j$ at year $t$, and diagonal entries $E^{(t)}_{ii}$ also include individuals who resided in area $i$ at year $t-1$ but died or emigrated. We process the Census datasets so they reflect the same population, by accounting for changes in the populations at year $t$ due to births, deaths, and international migration. Specifically, when using Census populations from year $t$ to scale entries of $E^{(t)}$, we want to remove from each Census area population the natural increase in population (i.e., births minus deaths) and the net international migration (i.e., immigrants minus emigrants); when using Census flows from year $t-1$ to year $t$, we want to add counts of deaths and emigrants back into the diagonal as non-movers. We adjust the Census constraints individually for every matrix $E^{(t)}$.
To do this adjustment, we make use of the PEP components of population change and ACS international immigration estimates. PEP releases estimates for the number of births, deaths, net international migration, and net domestic migration yearly at both a county and a state level [55]. To build yearly population estimates, PEP directly adds to the previous year’s estimates the natural increase, net international migration, and net domestic migration [^6] PEP produces estimates that are geographically consistent: county-level estimates must aggregate precisely to state-level estimates, which must aggregate to nation-level estimates [55]. To achieve such precision, they also estimate small residuals at the county and state levels. We ignore these values.. Whereas PEP estimates only report the net international migration, the ACS 1-year state-to-state flows also estimate the yearly number of immigrants per state (average CV around 12% across all years), which allows us to estimate the number of emigrants per state.
Appendix C Harmonizing Infutor and Census Data
Having processed the Infutor and Census data, the final step in producing MIGRATE is to harmonize both data sources. Our harmonization process consists of scaling selected entries of the $E^{(t)}$ to match non-zero entries of Census datasets. We do not scale any entries to zero because the Census zeroes are noisily estimated, and subsequent multiplicative re-scalings cannot correct erroneous zero scalings. We detail below, in order, the scalings we conduct.
C.1 Harmonizing with CBG population data
First, we harmonize $E^{(t)}$ with CBG populations from the 5-year ACS estimates and the 2010 Census. We scale each row of each matrix $E^{(t)}$ such that the row sums — i.e., the CBG populations — are consistent with the constraints implied by the ACS 5-year CBG populations and the 2010 Census. For example, when we take the average sum of each row across the five matrices $\{E^{(2011)},E^{(2012)},E^{(2013)},E^{(2014)},E^{(2015)}\}$, this should be equivalent to the population for the corresponding CBG in the 2011-2015 ACS. Imposing these constraints corresponds to solving a linear system for each row using non-linear least squares (to impose the constraint that populations cannot be negative).
C.2 Harmonizing with yearly state non-movers data
Next, we harmonize $E^{(t)}$ with the counts of movers and non-movers by state from 1-year ACS data. To do this, we aggregate the off-diagonal entries of columns corresponding to each state and scale them to match the count of movers by state. We also aggregate the diagonal entries of columns corresponding to each state and scale them to match the count of non-movers by state. These scalings (which treat the population remaining within each CBG as equivalent to the population of non-movers) assume that very few people move to another location within the same CBG, an assumption substantiated by previous research. [^7] Research has shown that median distance people moved in the US was about 10 to 15 miles in the 2010-19 period, which is far more than the average 2010 CBG radius (1.7 miles) and close to the 98-th percentile of radii [56].
C.3 Harmonizing with yearly state-to-state flow data
We then harmonize $E^{(t)}$ with all state-to-state flows from 1-year ACS data. To do this, we aggregate the entries of the matrix $E^{(t)}$ according to the origin-destination state pair they belong to. For example, all flows from CBGs within Delaware to CBGs within Nevada are aggregated and scaled so that they sum to the total flow from Delaware to Nevada in ACS data. We do not scale to match zero flows.
C.4 Harmonizing with yearly county population data
We finally harmonize $E^{(t)}$ with county populations from 1-year PEP data. We choose to match $E^{(t)}$ to PEP populations last due to their superior reliability. In addition, PEP estimates are used internally by the Census Bureau as controls for many other datasets [57], and ensuring MIGRATE estimates agree with PEP can lead to better downstream agreement in other datasets.
To match to PEP county populations, we apply an IPF-based algorithm, alternating between two steps until convergence: (1) we aggregate the entries of $E^{(t)}$ by columns corresponding to each county and scale them to match the PEP county population at year $t$ and (2) we aggregate the entries of $E^{(t)}$ by rows corresponding to each county and scale them to match the PEP county population at year $t-1$. Our algorithm runs for $3,000$ iterations, which we verify is sufficient for convergence. Specifically, at the end of 3,000 iterations, the total update to the matrix $E^{(t)}$—that is, the $L^{1}$ distance between the resulting matrices in two subsequent iterations—falls to the order of $10^{-7}$.
Appendix D Validations
Here we provide additional details on MIGRATE validations. §D.1 provides additional validation metrics that complement Figure 2; §D.2 presents additional details on our held-out data validations; §D.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §D.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
D.1 Additional validation metrics
Figure 6: Detailed flow validations of MIGRATE. To complement the findings shown in Figure 2, we present validations with respect to area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We show (a-c) Pearson correlations of estimates and selected ACS flows as well as (d-f) how much MIGRATE reduces RMSE compared to Infutor.
Figure 6 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We report these metrics at both the state and county level; state-to-state flows were used in the estimation procedure, county-to-county flows were held out. For all datasets, MIGRATE estimates are highly correlated with Census estimates ($\rho$ above $0.957$ in all cases) and reduce RMSE relative to the Infutor data.
D.2 Validating the estimation procedure using held-out data
MIGRATE CBG Held-Out State Held-Out
Population counts State 1.000 1.000 1.000
County 1.000 1.000 1.000
Tract 0.997 0.888 0.998
CBG 0.996 0.856 0.997
Population flows State 1.000 1.000 1.000
County 1.000 1.000 1.000
Population flows (area movers only) State 0.998 0.998 0.895
County 0.957 0.956 0.949
Population flows (area stayers only) State 1.000 1.000 1.000
County 1.000 1.000 1.000
In-migration State 0.998 0.998 0.885
County 0.984 0.983 0.963
In-migration rate State 0.987 0.987 0.663
County 0.715 0.705 0.629
(a) Pearson correlation between Census data and our estimates with CBG and state data held out. Pearson correlations with MIGRATE, which is estimated using all data sources, are presented for comparison.
MIGRATE CBG Held-Out State Held-Out
Population counts State 100.00 100.00 100.00
County 100.00 100.00 100.00
Tract 85.88 13.46 88.13
CBG 83.60 9.08 85.09
Population flows State 89.46 89.40 80.87
County 90.65 90.70 75.24
Population flows (area movers only) State 87.46 87.06 4.08
County 42.34 41.30 4.65
Population flows (area stayers only) State 89.46 89.40 81.03
County 91.01 91.08 75.59
In-migration State 90.93 91.11 7.22
County 68.92 68.19 6.55
In-migration rate State 87.30 87.33 4.89
County 51.75 50.91 0.73
(b) Reductions in RMSE. We compute RMSE between (1) our estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using our estimates. Positive values indicate improvements. Values are average reductions across all years with overlapping ground-truth data releases.
Table 4: Validation of estimates in held-out data settings. The full model validations are repeated for comparison. For in-migration rates, metrics are weighted by populations to account for disparate county and state sizes.
We verify that our estimation procedure yields highly correlated estimates with each Census data source even when it is removed from the datasets used for estimation. For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data. Specifically, we conduct our full battery of validations on two additional sets of migration estimates: “CBG held-out”, which is estimated without the CBG-level population data (i.e. $E^{(t)}$ was not scaled to match ACS 5-year CBG populations); and “State held-out”, which is estimated without the state-level flow data (i.e. $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates). Table 4 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor. Values for the original MIGRATE estimates, reported in Figure 2 and Figure 6, are reproduced for easy comparison.
The CBG held-out and State held-out estimates remain highly correlated with all Census datasets ($\rho$ above $0.629$); they also reduce RMSE relative to Infutor. Unsurprisingly, both the correlations and error reductions are smaller than those achieved by our full MIGRATE estimates, which make use of all sources of Census data. Together, these results illustrate both that (a) our estimation pipeline generalizes well to held-out data and that (b) all datasets used in the training help improve concordance with Census data, justifying their inclusion.
D.3 Related work documenting biases in the Infutor data
Many authors have assessed bias in the Infutor dataset, generally focusing on the data subsets or demographic characteristics most relevant to their own analyses. To our knowledge, the analysis we present in Figure 3 constitutes the first comprehensive documentation of biases in the Infutor dataset across (a) a long period of time and (b) the entire United States. We include here a non-exhaustive review of works with documentation of Infutor biases.
As we report in Table 2, Infutor undercounts the national US population. This is consistent with prior work; a similar finding is reported by Qian and Tan [19], Bernstein et al. [17], and Diamond et al. [15] among others. Phillips [16] additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers. A contrasting finding, reported by Diamond et al. [4], is that Infutor might overcount the population according to the 2000 Census in San Francisco. Such a finding might be driven by specific spatiotemporal idiosyncrasies of the San Francisco data in 2000, which is not included in the subset we analyze; however, the bulk of the literature, and all the literature which analyzes the same subset of the data we do, is consistent with our finding that Infutor undercounts the data. Diamond et al. [4] also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
In Figure 3, we also assess whether under- or over-represents demographic subpopulations. (Prior to conducting this assessment, we scale the Infutor population to match the overall population, so any discrepancies are not caused merely by overall undercounting.) Ramiller et al. [20] similarly study representation biases of Infutor in King County, Washington, during the 2015-2019 5-year period. They find that Infutor undercounts the county population while overrepresenting White, older, home-owning, and high-income populations. Later, they regress population errors in tract-level demographics to find that Infutor records are overrepresented in areas with higher White and high-income populations among other characteristics. We confirm most of their findings in this limited spatiotemporal scale at national level. However, the authors use the NARC3 dataset—a subset of Infutor’s address records that contains detailed demographic information—and thus, beyond the spatiotemporal scope, our analyses might be substantially different.
D.4 Assessment of additional demographic biases
Figure 7: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
We documented biases in the Infutor dataset geographically (Figure 3a) and on the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c). Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure 7). We find that Infutor displays negligible bias on the basis of sex. However, Infutor underrepresents populations with less formal education while overrepresenting populations with more formal education. Populations living in households with no children are overrepresented while those in households with children are underrepresented. This finding is consistent with the Infutor data documentation, which describes a lack of records for people under the age of 18 (therefore, children will be undercounted). MIGRATE essentially eliminates biases along all these demographic dimensions.
Appendix E Case Study: Flows from New York City and Income
Figure 8: Average median income of destination CBGs and counties for New York City out-movers in the 2010-19 period. The large map uses MIGRATE estimates and 5-year ACS CBG data. The inset map shows the corresponding distribution using only publicly available migration data from 5-year ACS county-to-county estimates.
We illustrate the value of MIGRATE estimates in an additional case study: contrasting MIGRATE with publicly available county-to-county migration data when analyzing migration stratified by a demographic or socioeconomic feature. Figure 8 plots the average median income of the CBG to which a person moves according to their previous CBG of residence. To compute this quantity for a given CBG $i$, we average the median income of all destination CBGs $j$ for out-movers of CBG $i$ weighted by the number of out-movers from $i$ to $j$. We perform an analogous computation for the county plot in the inset.
Comparing these results to the data available at county level reveals striking within-county variation. For example, the median income of the county to which people from Manhattan moved was \$$\$60,502$—lower than that of the other four counties in New York City. MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest median income among all New York City CBGs — above \$$\$100,000$ — while movers from other Manhattan CBGs moved to areas with some of the lowest median incomes. Intra-county socioeconomic differences factor heavily in out-migration: Manhattan out-movers with the highest-income destinations tended to reside in the central and lower parts of the county (which tend to have higher incomes), whereas those with lower-income destinations came mostly from northern CBGs (which tend to have lower incomes). Similar within-county dynamics happen in the other four counties. Overall, this analysis illustrates the important demographic heterogeneity in migration which is concealed at the county level.
Appendix F Privacy Protections
In this section we detail the measures we took to preserve the privacy of individuals in the MIGRATE data release. Analysis of the Infutor dataset was determined to be not human subjects research by the Cornell Institutional Review Board (IRB #0145225).
First, in contrast to the original Infutor data, the data we release is aggregated: it contains only estimated migration between CBG-CBG pairs, and no information about individuals. This level of data aggregation is consistent with other data sources which have been made available to researchers. For example, mobility data often contains the aggregate counts of individuals moving from a given CBG to a given point of interest on a given day [58, 59, 60, 61, 62].
Second, the matrices we release are probabilistic and rescaled: in particular, the fact that a particular entry of the CBG-CBG migration matrix is non-zero does not guarantee that a given number of individuals moved between those CBGs (or indeed, that any did at all). This is because individuals are mapped probabilistically to CBGs (e.g., if we have only a ZIP code for an individual, they are mapped probabilistically to CBGs within that ZIP code) and the migration matrices are rescaled multiple times.
Third, we release the data only to researchers for non-profit use who agree to a data use agreement pledging not to re-identify individuals in the data, following manual review of their application and research project.
Finally, we compute the average number of destinations (unique CBGs) for out-movers from each CBG. The purpose of the analysis is to assess how difficult it would be to track a given individual if one knew they lived in a given CBG and then moved. Across CBGs and years, the mean number of destinations for outmovers is $272.2$ (median $234$). We redact from the public data release migration information for the very small proportion of CBGs ($0.558\%$) for which 90% of out-movers travel to 10 or fewer CBGs. All analyses in the paper were conducted with the full data.

### v2

@@
 However, migration datasets within the United States remain lacking: publicly available Census data are neither spatially nor temporally granular, and proprietary data have higher resolution but demographic and other biases.
-To address these limitations, we develop a scalable iterative-proportional-fitting based method which reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data.
-We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 which captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data.
+To address these limitations, we develop a scalable iterative-proportional-fitting based method that reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data.
+We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 that captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data.
 These estimates are highly correlated with external ground-truth datasets, and improve accuracy and reduce bias relative to raw proprietary data.
-We publicly release MIGRATE estimates and provide a case study illustrating how they reveal granular patterns of migration in response to California wildfires.
+We use MIGRATE to analyze both national and local migration patterns.
+Nationally, we document temporal and demographic variation in homophily, upward mobility, and moving distance: for example, we find that people are increasingly likely to move to top-income-quartile CBGs and identify racial disparities in upward mobility.
+We also show that MIGRATE can illuminate important local migration patterns, including out-migration in response to California wildfires, that are invisible in coarser previous datasets.
+We publicly release MIGRATE to provide a resource for migration research in the social, environmental, and health sciences.
 1 Introduction
 Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena.
-Migration patterns illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; effects of the COVID-19 pandemic [8]; and political polarization [9].
-But migration data within the United States have serious limitations.
-The most fine-grained freely available datasets track migration at the county level.
+Migration data illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; consequences of conflicts [8]; effects of the COVID-19 pandemic [9]; and political polarization [10].
+But migration datasets within the United States have serious limitations.
+The most fine-grained freely available datasets track migration at the county level [11].
 While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level.
 For example, research on flood-risk-induced migration uses models at the much more granular Census block level [12], arguing that this spatial granularity is necessary due to the highly localized nature of flood risk [13].
-Research on neighborhoods which are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [12].
+Research on neighborhoods that are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [14].
 Movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 [15].
@@
 For example, Infutor and other consumer record datasets have been shown to overrepresent higher-income and majority-group populations in some settings [22].
-To address these limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates which combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data.
-To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm which repeatedly scales a noisy matrix to reconcile it with constraints [21], to reconcile the raw Infutor migration matrix with the more reliable Census constraints.
+To address the limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates that combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data.
+To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm that repeatedly scales a noisy matrix to reconcile it with constraints [23, 24, 25], to reconcile the raw Infutor migration matrix with the more reliable Census constraints.
 We accommodate uncertainty in address mapping and incorporate data on births, deaths, and immigration.
 The output of our method is a set of yearly inferred migration matrices at the Census block group (CBG) level from 2010-2019 covering the 50 states and the District of Columbia.
-Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows which are currently available on the 5-year level, and 18 million times more granular than the state-state flows which are currently available on the 1-year level.
+Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows publicly available on the 5-year level, and 18 million times more granular than the state-state flows publicly available on the 1-year level.
+Figure 1: MIGRATE estimates.
+We estimate annual migration flows between all pairs of Census block groups (CBGs) from 2010-2019.
+(a) Average MIGRATE estimates of out-migration rates across the entire United States.
+(b-c) MIGRATE estimates of out-migration rates within New York City.
+MIGRATE estimates reveal granular spatial patterns invisible in publicly available county-to-county data (inset plot (b)).
+Out-migration rates for CBGs with fewer than 100 people are omitted.
 We comprehensively validate MIGRATE by comparing it to external data sources.
@@
 We document systematic demographic biases in the Infutor data, which overcounts rural, older, white, and home-owning populations, and show that MIGRATE almost entirely eliminates these biases.
-To illustrate the benefits of granular migration data, we provide a case study using our estimates to study migration in response to California wildfires, showing that MIGRATE reveals dramatic increases in out-migration which county-level data obscures.
+We then use MIGRATE to analyze both national and local migration.
+Nationally, we reveal both temporal and demographic variation in migration homophily, upward mobility, and moving distance.
+We find, for example, that people are increasingly likely to move to top-income-quartile CBGs, but also provide evidence of racial disparities: movers from plurality Black CBGs are less likely, and movers from plurality Asian CBGs more likely, to move to top-income-quartile CBGs, and these disparities persist even when controlling for income of origin CBG.
+Locally, we show that MIGRATE can illuminate important migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data.
+Collectively, these analyses illustrate how MIGRATE reveals important migration patterns at both the national and local levels.
 To provide a foundation for more precise migration research in the social, environmental, and health sciences, we release MIGRATE for other researchers at migrate.tech.cornell.edu.
-2 Creating MIGRATE
+2 Results
+2.1 Creating MIGRATE
 Here we provide a high-level summary of our method for inferring fine-grained migration matrices by harmonizing Census and Infutor data.
-In the Appendix, we provide full details of processing the Infutor dataset (Appendix A), processing Census data (Appendix B), and harmonizing the two data sources (Appendix C).
+In the Methods Section, we provide full details of processing the Infutor dataset (§M.1), processing Census data (§M.2), and harmonizing both data sources (§M.3).
 Preprocessing the raw Infutor dataset
@@
 We use this information to compute the number of people moving between each pair of geographic areas.
-In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior.
-We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Appendix Table 2.
+In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior [^1] We correct populations to account for birth, deaths, and international migration.
+Full details in the Methods Section (§M.2)..
+We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Table M1.
 Individuals with active records (any time in 2010-2019) 374,217,253
@@
 We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period.
-Comparing the average yearly number of active individuals in the Infutor data to the Census population shows that Infutor under-counts the population.
-(Table 2 provides the number of active individuals in each year from 2010-2019).
+Comparing the average yearly number of active individuals in the Infutor data to the Census population (second row) shows that Infutor under-counts the population.
+(Table M1 provides the number of active individuals in each year from 2010-2019).
 Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
@@
 We are able to map 99.21% of all addresses in Infutor to CBGs.
-We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in Appendix A.
+We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in the Methods Section (§M.1).
 Harmonizing the Infutor and Census datasets
@@
 We rescale to different Census datasets in sequence, rather than simultaneously, because the datasets are not totally consistent with each other.
-Figure 1: MIGRATE estimates.
-We estimate annual migration flows between all pairs of Census block groups (CBGs) from 2010-2019.
-(a) Average MIGRATE estimates of out-migration rates across the entire United States.
-(b-c) MIGRATE estimates of out-migration rates within New York City.
-MIGRATE estimates reveal granular spatial patterns invisible in publicly available county-to-county data (inset plot (b)).
-Out-migration rates for CBGs with fewer than 100 people are omitted.
 Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations; then the 1-year counts of movers and non-movers by state; then the 1-year state-to-state flows; and finally the 1-year county populations.
 The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them: while CBG populations and state flows are estimated by the American Community Survey (ACS) and often contain large sampling errors, county populations are estimated by the Population Estimates Program (PEP) using more precise large-scale administrative datasets.
-We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (see Appendix D.2).
-To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [21, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges (Appendix C.4).
+We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (Appendix A.2).
+To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [23, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges.
 The resulting migration matrices constitute our MIGRATE estimates.
-Appendix C provides additional methodological details on these scalings.
+The Methods Section (§M.3) provides additional methodological details on these scalings.
 Naive implementations of our harmonization algorithm impose prohibitive computation times due to the size of the matrices involved.
@@
 This granularity often reveals important patterns which are invisible at the county-county level, as we illustrate by zooming in on New York City and comparing the county-level rates obtained from ACS flows (Figure 1b) to the CBG-level out-migration rates inferred by MIGRATE (Figure 1c).
-In Appendix E, we include an additional analysis showing how the socioeconomic stratification of movers’ destinations changes depending on whether we look at county-level estimates or MIGRATE estimates: residents from lower-income CBGs in higher-income counties tend to move to other lower-income CBGs, but this trend is obscured at the county level.
-3 Validating MIGRATE
+2.2 Validating MIGRATE
 We conduct extensive validations of the MIGRATE estimates.
@@
 Figure 2: Validating the MIGRATE estimates.
-MIGRATE estimates are highly correlated with (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population).
+(a - c): MIGRATE estimates (y-axis) are highly correlated with Census data (x-axis), including (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population).
 For in-migration rates, correlations are weighted by state or county population, and points are sized by population.
@@
 MIGRATE estimates are also highly correlated with state-to-state flows ($\rho=1$); this high correlation is driven in part by variation in state populations, as well as by the fact that most people do not move.
-Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix D.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
+Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix A.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
 Importantly, we further verify that our estimation procedure yields highly correlated estimates even with datasets which are not used in our estimation procedure.
@@
 Specifically, we compare our MIGRATE estimates to ACS 5-year county-to-county flows, which are not incorporated in our estimation, and find they are highly correlated with our estimates ($\rho=1$).
-This high correlation is driven in part by the fact that most people do not move; hence, in Appendix D.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$).
+This high correlation is driven in part by the fact that most people do not move; hence, in Appendix A.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$).
 Another contributor to the high correlation in county-county flows is variation in county size, so we separately compute correlations with county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) which normalize for county size.
@@
 As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation: for example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census block group populations).
-See Appendix D.2 for full results.
+See Appendix A.2 for full results.
 Overall, these validations demonstrate that MIGRATE estimates are highly correlated with external ground-truth Census data.
@@
 Nor is it due merely to the fact that most people do not move (since it remains true when examining correlations for movers specifically).
-Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets which are not used in our estimation procedure.
+Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets that are not used in our estimation procedure.
 MIGRATE estimates reduce error relative to raw Infutor data
@@
 MIGRATE estimates reduce error in state-to-state flows by $89.5\%$ and error in state in-migration rates by $83.3\%$.
-We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix D.1.
+We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix A.1.
 We again confirm that we still see these reductions in error on held-out data which is not used in our estimation procedure.
-Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$.
-Neither of these datasets was used in producing MIGRATE estimates.
-We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix D.2).
+Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$; neither of these datasets was used in producing MIGRATE estimates.
+We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix A.2).
 Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data.
@@
 Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates.
-Infutor data displays biases which MIGRATE estimates greatly reduce.
+Infutor data displays biases that MIGRATE estimates greatly reduce.
 (a) Average errors in county populations in Infutor data relative to Census data; orange denotes counties where Infutor underrepresents the population, and purple denotes counties where Infutor overrepresents it.
@@
 Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups.
-These biases are consistent with biases found in previous research [20], as we discuss in detail in Appendix D.3, and could propagate into biases in downstream analyses of inequality and other topics.
+These biases are consistent with biases found in previous research [22], as we discuss in detail in Appendix A.3, and could propagate into biases in downstream analyses of inequality and other topics.
 MIGRATE estimates reduce biases in Infutor data.
@@
 The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases.
-We present analyses for additional demographic groups, finding similar results, in Appendix D.4.
-4 Case study: migration in response to California wildfires
+We present analyses for additional demographic groups, finding similar results, in Appendix A.4.
+2.3 Analysis of national migration patterns
+Figure 4: National migration statistics.
+(a) Flows between ten types of CBGs — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile.
+Rows correspond to the origin CBG, and columns to the destination CBG; for example, the top left entry indicates that 90% of movers from plurality white CBGs move to plurality white CBGs.
+The final two rows report the proportion of all movers moving to CBGs of each type, and the population share living in CBGs of each type.
+We report averages across all years.
+(b) Probability of moving to a higher-median-income CBG, conditional on income decile of origin CBG, and plurality race of origin CBG.
+(c) Distance moved stratified by CBG type.
+We use MIGRATE to analyze national patterns of migration from 2010-2019.
+The spatial granularity of the data affords a unique opportunity to study demographic variation in migration patterns — for example, differences in migration between higher-income and lower-income CBGs — that may be obscured by county or state-level data, which show far less demographic variation.
+We hence divide CBGs into 10 (overlapping) categories — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile — and stratify the migration statistics we compute by these ten categories [^4] To determine the category of each CBG every year, we use the most recent ACS 5-year demographic estimates.
+For example, to classify CBGs into plurality races to study moves in the 2010-2011 period, we use the ACS 2006-2010 race and ethnicity estimates..
+Figure 4a plots flows between the ten categories — for example, the proportion of movers from top-income-quartile CBGs who move to top-income-quartile CBGs.
+This reveals substantial homophily in migration: out-movers from all ten CBG categories are more likely than movers as a whole to move to CBGs of the same category.
+However, the strength of this homophily varies across categories.
+Movers from plurality Black, Asian, and Hispanic CBGs exhibit strong homophily: they are 5.5$\times$, 14.6$\times$, and 4.3$\times$ likelier than movers as a whole to move to CBGs with the same plurality race group (Figure S3a reports relative rates for all groups).
+We also observe income homophily, particularly for movers within top and bottom quartile CBGs: movers from bottom-income-quartile CBGs are nearly twice as likely as movers as a whole to move to CBGs in the bottom income quartile (34% versus 18% for movers as a whole); movers from top-income-quartile CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs (53% versus 32%).
+In Figure S3b, we confirm that this homophily does not occur merely because many moves are local and demographics are spatially correlated: we also observe homophily when restricting to long-distance (out-of-county) moves.
+Figure 4a also shows that people are likelier to move to CBGs in higher income quartiles: 32% of moves are to top-income-quartile CBGs, while only 18% are to bottom-income-quartile CBGs.
+This trend becomes more pronounced over the decade we study (Figure S4d) and is only partially explained by the fact that top-income-quartile CBGs account for a larger share of the population (29%) than bottom-income-quartile CBGs (21%).
+There are also racial disparities: movers from plurality Asian CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs; movers from plurality Black CBGs are 2.0$\times$ likelier than movers as a whole to move to bottom-income-quartile CBGs.
+Figure 4b investigates whether these racial disparities persist when controlling for origin CBG median income: in particular, plotting the share of movers moving to a higher-income CBG when controlling for the mover’s origin CBG median income decile.
+This reveals that the probability of moving to a higher-income CBG varies substantially by race even conditional on origin CBG income: movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs.
+For example, movers from fifth-income-decile, plurality Asian CBGs have a 71% chance of moving to a higher-income CBG; movers from fifth-income-decile, plurality Black CBGs have a 53% chance.
+Figure S4 shows that the same racial disparities occur when controlling for CBG median income percentile, as opposed to decile; or when examining the probability of moving to a top- or bottom-income-quartile CBG, as opposed to moving to a higher-income CBG.
+Overall, we find robust and substantial racial disparities in income of destination CBG that persist even conditional on income of origin CBG.
+Finally, Figure 4c provides statistics on the distance of moves.
+37% of movers move less than 5 miles; 40% from 5-50 miles; and 23% more than 50 miles.
+Figure 4c also highlights demographic variation in these statistics, revealing that movers from plurality white CBGs, rural CBGs, and higher-income CBGs are likelier to move long distances (more than 50 miles).
+Appendix B.4 reports additional migration distance statistics stratified by geographic boundary (i.e., whether movers remain within the same tract, county, or state) and shows that migration distance increases over the decade we study.
+Overall, these results demonstrate that fine-grained migration data illuminates important demographic variation in homophily, upward mobility, and migration distance.
+2.4 Analysis of local migration patterns
 Figure 5: Migration in response to California wildfires.
@@
 (c) Out-migration rates in ACS 5-year county-to-county data remain relatively constant over time.
-We demonstrate the value of the MIGRATE estimates by using them to study migration in response to wildfires in California.
+In addition to using MIGRATE to analyze national migration trends, we use it to study local migration patterns: specifically, migration in response to wildfires in California.
 Natural disasters, including wildfires, are known drivers of human mobility [3, 29].
@@
 Looking at administrative address data from mail delivery after Camp Fire, for example, researchers found a much more significant uptick in short-term vacancy than in long-term vacancy [32].
-In contrast, CBGs which do not lie within the fire perimeters experience much lower out-migration rates.
+In contrast, CBGs that do not lie within the fire perimeters experience much lower out-migration rates.
 We systematically quantify these differences in Figure 5b, which compares CBGs within the fire perimeter to three sets of less-affected CBGs outside the perimeter: other CBGs neighboring the perimeter; other CBGs within the affected counties; and others within California as a whole.
@@
 (For this analysis, we compare to the relative change in the ACS 5-year dataset released the year following the fire from the ACS 5-year dataset the year before: for example, for the 2018 Camp Fire, we compare the 2015-2019 ACS estimates to the 2014-2018 ACS estimates.) The ACS 5-year population estimates, unlike MIGRATE estimates, also provide no insight into the destinations of out-movers.
-Collectively, these analyses showcase how MIGRATE estimates reveal substantial, policy-relevant responses to wildfires which traditional data sources conceal.
-Indeed, the MIGRATE estimates enable further analyses beyond those conducted here: for example, they can illuminate inequality in responses to wildfires by connecting out-migration patterns with CBG-level demographics.
+In Appendices C and D, we provide two additional analyses of local migration patterns that are impossible using only county-level data: we analyze socioeconomic variation in New York City out-mover destinations, and migration patterns for residents who live in public housing provided by the New York City Housing Authority as part of the city’s affordable housing policy.
+Collectively, these analyses showcase how MIGRATE estimates reveal important, policy-relevant local migration patterns that traditional data sources conceal.
 3 Discussion
@@
 Our estimates are approximately 4,600 times more spatially granular than publicly available county-to-county 5-year migration data, and 18 million times more spatially granular than state-to-state 1-year migration data.
-We produce these estimates by developing an IPF-based algorithm to harmonize granular but biased proprietary migration data with more reliable but less granular Census datasets.
-We show that MIGRATE estimates correlate highly with external Census ground-truth datasets and reduce error and demographic bias relative to the raw proprietary data.
-We discuss the measures taken to protect the privacy of all individuals in the dataset in Appendix F.
+MIGRATE estimates correlate highly with external Census ground-truth datasets and reduce error and demographic bias relative to raw proprietary data.
+We discuss the measures taken to protect the privacy of all individuals in the dataset in Appendix E.
 Our estimates are available at migrate.tech.cornell.edu.
+We use MIGRATE to analyze both local and national patterns of migration.
+We show that MIGRATE can reveal important local migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data.
+Nationally, we analyze demographic and temporal variation in migration patterns.
+We find that people tend to move to CBGs of the same type across dimensions of race, income, and rural/urban status — for example, movers from plurality Black, Asian, and Hispanic CBGs are 15.5$\times$, 14.6$\times$, and 4.3$\times$ likelier to move to CBGs with the same plurality race group, consistent with prior work documenting the persistence of racial segregation [51, 52, 53].
+We document demographic and temporal variation in moving distance: movers from plurality white, rural, and higher-income CBGs are likelier to travel long distances, and moving distance increases over the decade we study, consistent with prior work [15].
+An important finding from our national analysis is racial and temporal variation in upward mobility.
+Movers are likelier to move to top-income-quartile CBGs than bottom-income-quartile CBGs, a trend that becomes more pronounced over the decade we study.
+We find large racial disparities in the likelihood of moving to higher-income CBGs, or top-income-quartile CBGs, even when controlling for income of origin CBG: movers from plurality Asian CBGs are likelier to move to a higher-income CBG, and movers from plurality Black CBGs are less likely.
+These findings are consistent with prior work documenting racial disparities in neighborhood attainment [54, 55, 56, 57] and generalize these findings to larger and more recent samples and additional racial groups.
+Overall, we document demographic and temporal variation in national migration trends using a recent, large-scale, and highly granular dataset.
 While we validate MIGRATE estimates comprehensively against external data sources, practitioners making use of the data should be mindful of two limitations.
@@
 For example, while we correct biases in CBG populations, we cannot correct sub-CBG-level population biases, or biases in CBG-CBG flows.
-Second, while we show that MIGRATE estimates are highly correlated with external Census datasets (including held-out datasets which are not included in producing our estimates), these validations have imperfections.
-As we further describe in Appendix B.2, Census datasets themselves have biases; can possess significant margins of error, particularly at fine spatial scales; and are not perfectly consistent with each other.
+Second, while we show that MIGRATE estimates are highly correlated with external Census datasets (including held-out datasets that are not used to produce our estimates), these validations have imperfections.
+As we further describe in the Methods Section (§M.2), Census datasets themselves have biases; can possess significant margins of error, particularly at fine spatial scales; and are not perfectly consistent with each other.
 Another challenge in validation is the lack of a ground-truth CBG-CBG migration matrix against which to validate; publicly available Census datasets do not afford this level of granularity, and non-Census data sources (e.g., voter files) have significant biases and measure substantively different populations than the one we seek to model [58, 59].
-Overall, MIGRATE should be viewed as a new data source which like all data possesses limitations but nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
+Overall, MIGRATE should be viewed as a new data source which like all data possesses limitations but which nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
 We hope these improvements will facilitate a wide range of further migration-related analyses.
-As our wildfire case study illustrates, MIGRATE reveals patterns which remain invisible in publicly available migration data.
-The fine spatial scale of the data also enables analyses of inequality in migration, by linking migration patterns to CBG-level demographics.
-Our work also highlights methodological directions for improving estimation of fine-grained migration patterns, including providing principled margins of error on estimates and incorporating additional data sets to further improve estimation and validation.
-More broadly, we hope MIGRATE will facilitate more precise study of many migration-related phenomena across the social, health, and environmental sciences.
-Appendix A Processing Infutor Data
+As our analyses illustrate, MIGRATE reveals patterns which remain invisible in publicly available, county-level migration data.
+We hope MIGRATE will facilitate more precise study of many migration-related phenomena across the social, health, and environmental sciences.
+Methods
+M.1 Processing Infutor Data
 Infutor provides data in tabular form.
@@
 We describe the steps taken to (1) identify and clean Infutor data records within the scope of MIGRATE, (2) process the Infutor data into matrices documenting yearly flows between address pairs, and (3) map these address-level matrices to the CBG-level matrices $E^{(t)}$ mentioned in the main text.
-Table 2 summarizes the raw and processed data.
-A.1 Cleaning address histories
+Table M1 summarizes the raw and processed data.
+Cleaning address histories
 We first create a sequence of monthly addresses for each individual in Infutor: i.e., the address at which they are living each month.
@@
 At the end of this process, each individual is associated with a monthly probability distribution over addresses for each month in their interval of activity.
-A.2 Processing address histories into annual address-address migration matrices
+Processing address histories into annual address-address migration matrices
 When then aggregate the monthly address histories for each individual to the annual level in a way which is consistent with the ACS interview process.
@@
 These matrices will be further aggregated based on the location of each address.
-A.3 Processing address-address migration matrices into CBG-CBG migration matrices
-Figure 5: Flowchart detailing the process of mapping addresses to CBGs.
+Processing address-address migration matrices into CBG-CBG migration matrices
+Fig.
+M1: Flowchart detailing the process of mapping addresses to CBGs.
 We start with all Infutor addresses and classify them into five address types.
@@
 We use the 2010 Census block group boundaries for all addresses to maintain geographic consistency.
-Figure 5 details this process.
+Figure M1 details this process.
 Our address mapping pipeline is divided in two parts: mapping addresses via geocoding, and mapping addresses with a ZIP code-to-CBG crosswalk.
@@
 We then distribute the probability to each Census block group within the tracts, again weighting by their relative population shares.
-Success rates for each of these steps are high and detailed in Figure 5.
+Success rates for each of these steps are high and detailed in Figure M1.
 We use the results of this mapping process to create an auxiliary matrix $\mathcal{G}$ of dimensions $N_{\text{ADDRESSES}}\times N_{\text{CBGs}}$.
@@
 The second term in this equation maps stayers to the appropriate diagonal entries of the migration matrix: for example, a precisely mapped stayer in CBG $i$ will increment $E^{(t)}_{ii}$ by 1, and someone who stays within ZIP code $a$ will distribute mass (summing to 1) evenly across the diagonal entries corresponding to CBGs within that ZIP code.
-Table 2 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline.
+Table M1 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline.
 The number of individuals with active records in the data represents the number of individuals with an interval of activity containing each year at the start of the data processing; the number of CBG moves corresponds to the sum of off-diagonal entries of each matrix $E^{(t+1)}$ and measures the “effective size” of each year’s data: how many movers we actually account for at the beginning of the estimation procedure.
-Year (1) Active records (2) US Population (3) Effective Dates (4) Processed CBG Moves
-2010 269,228,776 309,327,143 48,225,002 15,180,982
-2011 260,064,921 311,583,481 28,196,562 11,793,806
-2012 259,451,204 313,877,662 32,009,822 12,544,199
-2013 252,933,953 316,059,947 27,386,564 13,829,056
-2014 256,453,630 318,386,329 37,386,684 13,071,070
-2015 241,295,957 320,738,994 30,777,204 14,022,249
-2016 242,298,262 323,071,755 38,085,337 13,605,804
-2017 193,479,001 325,122,128 34,599,168 12,684,184
-2018 177,629,604 326,838,199 33,183,982 13,505,890
-Table 2: Yearly Breakdown of Data Summaries.
+Year (1) Active records (2) US Population (3) Processed CBG Moves
+2010 269,228,776 309,327,143 15,180,982
+2011 260,064,921 311,583,481 11,793,806
+2012 259,451,204 313,877,662 12,544,199
+2013 252,933,953 316,059,947 13,829,056
+2014 256,453,630 318,386,329 13,071,070
+2015 241,295,957 320,738,994 14,022,249
+2016 242,298,262 323,071,755 13,605,804
+2017 193,479,001 325,122,128 12,684,184
+2018 177,629,604 326,838,199 13,505,890
+Table M1: Yearly Breakdown of Data Summaries.
 Population and addresses accounted by the Infutor and Census datasets during the analysis period (2010-2019).
@@
 Comparison of these values to (2) the Census population shows that Infutor under-counts the population at every given year.
-(3) The number of effective dates per year is naturally reduced as we process the data into (4) CBG-to-CBG moves: an individual with two addresses will have at most one move, some addresses do not constitute residences, some individuals may have a single address in the year thus not move, etc.
-Appendix B Processing Census Data
+(3) The number of CBG-to-CBG moves parsed in our dataset.
+Note that an individual with two addresses will have at most one move, some addresses do not constitute residences, some individuals may have a single address in the year thus not move, etc.
+M.2 Processing Census Data
 The U.S.
@@
 We use a series of steps to clean, select, and process Census datasets.
-We describe our procedures for reconciling Census geographies (§B.1), selecting which Census datasets to use as constraints (§B.2), and how we process the Census data to deal consistently with births, deaths, emigration, and immigration (§B.3).
-B.1 Reconciling Census geographies
+We describe our procedures for reconciling Census geographies, selecting which Census datasets to use as constraints, and how we process the Census data to deal consistently with births, deaths, emigration, and immigration.
+Reconciling Census geographies
 The United States is hierarchically divided into states, counties, Census tracts, Census block groups, and Census blocks, with each level subdividing its parent geography.
@@
 After these adjustments, MIGRATE considers a universe of $217,740$ Census block groups grouped into $3,142$ counties.
-B.2 Selecting Census datasets
+Selecting Census datasets
 2010 2011 2012 2013 2014 2015 2016 2017 2018 2019
@@
 State In-Migration Rate - 4.95 4.68 4.79 4.65 4.70 4.76 4.72 4.85 5.02
-Table 3: ACS Coefficients of Variation.
+Table M2: ACS Coefficients of Variation.
 Average coefficient of variation (in %) of the non-zero estimates in each dataset.
@@
 We quantify the sampling error by the coefficient of variation (CV): the ratio of the standard deviation to the estimate value.
-Table 3 reports the mean coefficient of variation in ACS estimates.
+Table M2 reports the mean coefficient of variation in ACS estimates.
 In general, the number of non-movers has very low CVs (mean below 2% for counties and below 1% for states); population estimates have considerably lower CVs than flow estimates; state-to-state flows have relatively high CVs (mean around 45% every year), and county-to-county flows have extremely high CVs (mean almost 90%).
@@
 Based on these analyses, we harmonize the entries of each yearly matrix $E^{(t)}$ with the following data: (1) CBG populations (both from 2010 Census and 5-year ACS); (2) ACS 1-year counts of movers and non-movers by state; (3) ACS 1-year state-to-state flows; and (4) PEP 1-year county populations.
-B.3 Accounting for natural population increase and international migration
+Accounting for natural population increase and international migration
 Using Census datasets to constrain MIGRATE estimates requires ensuring that the Census datasets reflect the same population accounted for in the Infutor matrix $E^{(t)}$.
@@
 Whereas PEP estimates only report the net international migration, the ACS 1-year state-to-state flows also estimate the yearly number of immigrants per state (average CV around 12% across all years), which allows us to estimate the number of emigrants per state.
-Appendix C Harmonizing Infutor and Census Data
+M.3 Harmonizing Infutor and Census Data
 Having processed the Infutor and Census data, the final step in producing MIGRATE is to harmonize both data sources.
@@
 We detail below, in order, the scalings we conduct.
-C.1 Harmonizing with CBG population data
+Harmonizing with CBG population data
 First, we harmonize $E^{(t)}$ with CBG populations from the 5-year ACS estimates and the 2010 Census.
@@
 Imposing these constraints corresponds to solving a linear system for each row using non-linear least squares (to impose the constraint that populations cannot be negative).
-C.2 Harmonizing with yearly state non-movers data
+Harmonizing with yearly state non-movers data
 Next, we harmonize $E^{(t)}$ with the counts of movers and non-movers by state from 1-year ACS data.
@@
 [^9] Research has shown that median distance people moved in the US was about 10 to 15 miles in the 2010-19 period, which is far more than the average 2010 CBG radius (1.7 miles) and close to the 98-th percentile of radii [66].
-C.3 Harmonizing with yearly state-to-state flow data
+Harmonizing with yearly state-to-state flow data
 We then harmonize $E^{(t)}$ with all state-to-state flows from 1-year ACS data.
@@
 We do not scale to match zero flows.
-C.4 Harmonizing with yearly county population data
+Harmonizing with yearly county population data
 We finally harmonize $E^{(t)}$ with county populations from 1-year PEP data.
@@
 Specifically, at the end of 3,000 iterations, the total update to the matrix $E^{(t)}$—that is, the $L^{1}$ distance between the resulting matrices in two subsequent iterations—falls to the order of $10^{-7}$.
-Appendix D Validations
+[…]
+Supplementary information Supplementary Information is available detailing additional methods, validations, and findings.
+Declarations
+[…]
+Data availability MIGRATE is available upon request for non-profit research use at migrate.tech.cornell.edu.
+[…]
+Code availability Code to reproduce our research findings is available on GitHub: https://github.com/gsagostini/MIGRATE.
+Appendix A Additional details on validations
 Here we provide additional details on MIGRATE validations.
-§D.1 provides additional validation metrics that complement Figure 2; §D.2 presents additional details on our held-out data validations; §D.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §D.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
-D.1 Additional validation metrics
-Figure 6: Detailed flow validations of MIGRATE.
+§A.1 provides additional validation metrics that complement Figure 2; §A.2 presents additional details on our held-out data validations; §A.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §A.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
+A.1 Additional validation metrics
+Fig.
+S1: Detailed flow validations of MIGRATE.
 To complement the findings shown in Figure 2, we present validations with respect to area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e.
@@
 We show (a-c) Pearson correlations of estimates and selected ACS flows as well as (d-f) how much MIGRATE reduces RMSE compared to Infutor.
-Figure 6 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e.
+Figure S1 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e.
 the off-diagonal entries of the migration matrix), and raw in-migration (i.e.
@@
 For all datasets, MIGRATE estimates are highly correlated with Census estimates ($\rho$ above $0.957$ in all cases) and reduce RMSE relative to the Infutor data.
-D.2 Validating the estimation procedure using held-out data
+A.2 Validating the estimation procedure using held-out data
 MIGRATE CBG Held-Out State Held-Out
@@
 Values are average reductions across all years with overlapping ground-truth data releases.
-Table 4: Validation of estimates in held-out data settings.
+Table S1: Validation of estimates in held-out data settings.
 The full model validations are repeated for comparison.
@@
 $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates).
-Table 4 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor.
-Values for the original MIGRATE estimates, reported in Figure 2 and Figure 6, are reproduced for easy comparison.
+Table S1 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor.
+Values for the original MIGRATE estimates, reported in Figure 2 and Figure S1, are reproduced for easy comparison.
 The CBG held-out and State held-out estimates remain highly correlated with all Census datasets ($\rho$ above $0.629$); they also reduce RMSE relative to Infutor.
@@
 Together, these results illustrate both that (a) our estimation pipeline generalizes well to held-out data and that (b) all datasets used in the training help improve concordance with Census data, justifying their inclusion.
-D.3 Related work documenting biases in the Infutor data
+A.3 Related work documenting biases in the Infutor data
 Many authors have assessed bias in the Infutor dataset, generally focusing on the data subsets or demographic characteristics most relevant to their own analyses.
@@
 We include here a non-exhaustive review of works with documentation of Infutor biases.
-As we report in Table 2, Infutor undercounts the national US population.
+As we report in Table M1, Infutor undercounts the national US population.
 This is consistent with prior work; a similar finding is reported by Qian and Tan [21], Bernstein et al.
@@
 However, the authors use the NARC3 dataset—a subset of Infutor’s address records that contains detailed demographic information—and thus, beyond the spatiotemporal scope, our analyses might be substantially different.
-D.4 Assessment of additional demographic biases
-Figure 7: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
+A.4 Assessment of additional demographic biases
+Fig. S2: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
 We documented biases in the Infutor dataset geographically (Figure 3a) and on the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c).
-Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure 7).
+Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure S2).
 We find that Infutor displays negligible bias on the basis of sex.
@@
 MIGRATE essentially eliminates biases along all these demographic dimensions.
-Appendix E Case Study: Flows from New York City and Income
-Figure 8: Average median income of destination CBGs and counties for New York City out-movers in the 2010-19 period.
+Appendix B Additional analyses of national migration
+In this section we provide additional analyses of national migration trends discussed in Section 2.3.
+B.1 Out-migration rates by CBG type
+Out-migration rate (per 1,000) Out-movers
+Plurality Race White 132 31,425,400
+Black 128 3,823,677
+Asian 124 889,755
+Hispanic 118 5,289,710
+Urbanization Urban 135 35,481,142
+Rural 104 5,947,401
+Income Quartile Bottom 130 8,621,155
+2nd 130 9,944,217
+3rd 131 11,011,339
+Top 128 11,851,831
+Table S2: Out-migration by group.
+Out-migration rates (per 1,000 people) and raw counts of out-movers categorized by plurality race, urbanization level, and income quartile.
+We average the values across all years in the period.
+Table S2 reports the average out-migration rates for each of the ten CBG groups we analyze in the main text.
+The average out-migration rate in all CBGs is $13.0\%$, with $0.9\%$ standard deviation across years.
+Out-migration rates are lower in plurality Hispanic and rural CBGs; out-migration rates remain similar across income quartiles.
+B.2 Additional statistics on migration homophily
+Fig.
+S3: Additional national migration statistics.
+(a) Flows between the ten types of CBGs discussed in the main text relative to the share of all movers who moved to CBGs in the particular group.
+Each cell reports the ratio share of movers from origin CBG group moving to destination CBG groupshare of all movers moving to destination CBG $\frac{\text{share of movers from origin CBG group moving to destination CBG group}}{\text{share of all movers moving to destination CBG group}}$; for example, the top left value of $\times1$ reports the ratio of the share of movers from plurality white CBGs who move to plurality white CBGs (90%) to the share of movers from all CBGs who move to plurality white CBGs (80%).
+(b) Flows between the ten types of CBGs discussed in the main text when restricting to out-of-county movers demonstrate that homophily persists for long-distance moves.
+In Figure S3, we extend the analysis of national homophily patterns in migration documented in Figure 4a.
+Figure S3a shows the shares of movers moving between groups of CBGs relative to the share of all movers who move to that group of CBGs (that is, a version of Figure 4a where the value in each cell is divided by the corresponding value in the row ‘Movers from all CBGs’).
+Homophily in migration might occur because many moves are local (within county or within 5 miles) and demographics are spatially correlated.
+To investigate this possibility, we restrict our analysis to out-of-county moves in Figure S3b.
+Absolute racial homophily is less prominent—i.e.
+there are smaller shares in the diagonal of the race by race submatrix—yet still present: out-of-county movers from plurality Black, Asian, and Hispanic CBGs are 4.3$\times$, 5.5$\times$, and 3.1$\times$ likelier than out-of-county movers as a whole to move to CBGs with the same plurality race group (vs.
+5.5$\times$, 14.6$\times$, and 4.3$\times$ for all movers).
+B.3 Additional statistics on economic mobility in migration
+Fig.
+S4: Further national migration statistics on upwards mobility.
+Probability of moving to a (a) top-income quartile CBG, (b) bottom-income quartile CBG, and (c) higher-income CBG conditioned on current residence median income and plurality race.
+In (c), we classify current residence CBGs by income percentiles as opposed to deciles (as in the main text).
+Disparities remain robust.
+(d) Movers remain more likely to move to top-income-quartile CBGs than bottom-income-quartile CBGs throughout the decade, with the disparity increasing in magnitude over time.
+Figure S4 provides additional analyses supporting our finding of racial disparities in economic mobility (Figure 4b): movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs, even when controlling for income of origin CBG.
+We show that analogous racial disparities still emerge when looking at probability of moving to a CBG in the top income quartile (Figure S4a) or bottom income quartile (Figure S4b) as opposed to the probability of moving to a higher-income CBG (as in the main text).
+These disparities also emerge when controlling for origin CBG income percentile (Figure S4c), as opposed to decile, as in the main text.
+We also observe that the gap in upward mobility widens during the decade: the share of movers going to a top income quartile CBG increases from 31% to 33% while the share of movers going to a bottom income quartile CBG decreases from 19% to 17%; overall, the ratio of out-movers moving to top income quartile versus bottom income quartile CBGs increases from 1.64 to 1.95.
+B.4 Additional statistics on moving distance
+Fig. S5: Further national migration statistics on distance moved. (a) Moving distance by year. (b) Fraction of movers moving within tract, county, and state.
+Figure S5 provides additional statistics on moving distance, expanding the findings in Figure 4c.
+Figure S5a reports how moving distance changes over time, showing a general increase over the decade: the share of movers moving over 50 miles, for example, grows from about 22% to about 26%, and the average moving distance increases from 154 to 191 miles in the decade.
+Figure S5b reports moving distance stratified by geographic boundary (that is, the share of movers remaining within tract, county, and state).
+Findings are generally consistent with those reported in Figure 4c, and also with published ACS figures, which state that over half of moves happen within counties [15].
+Appendix C Case Study: destinations of New York City outmovers
+Fig.
+S6: Average income of destination CBGs and counties for New York City out-movers in the 2010-19 period.
 The large map uses MIGRATE estimates and 5-year ACS CBG data.
 The inset map shows the corresponding distribution using only publicly available migration data from 5-year ACS county-to-county estimates.
-We illustrate the value of MIGRATE estimates in an additional case study: contrasting MIGRATE with publicly available county-to-county migration data when analyzing migration stratified by a demographic or socioeconomic feature.
-Figure 8 plots the average median income of the CBG to which a person moves according to their previous CBG of residence.
-To compute this quantity for a given CBG $i$, we average the median income of all destination CBGs $j$ for out-movers of CBG $i$ weighted by the number of out-movers from $i$ to $j$.
+We provide a case study contrasting MIGRATE with publicly available county-to-county migration data when analyzing migration stratified by a demographic or socioeconomic feature.
+Figure S6 plots the average income of the CBG to which a person moves according to their previous CBG of residence.
+To compute this quantity for a given CBG $i$, we average the mean income of all destination CBGs $j$ for out-movers of CBG $i$ weighted by the number of out-movers from $i$ to $j$.
 We perform an analogous computation for the county plot in the inset.
 Comparing these results to the data available at county level reveals striking within-county variation.
-For example, the median income of the county to which people from Manhattan moved was \$$\$60,502$—lower than that of the other four counties in New York City.
-MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest median income among all New York City CBGs — above \$$\$100,000$ — while movers from other Manhattan CBGs moved to areas with some of the lowest median incomes.
+For example, the average income of the county to which people from Manhattan moved was \$$\$82,934$—lower than that of the other four counties in New York City.
+MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest average income among all New York City CBGs — above \$$\$175,000$ at times — while movers from other Manhattan CBGs moved to areas with some of the lowest average incomes.
 Intra-county socioeconomic differences factor heavily in out-migration: Manhattan out-movers with the highest-income destinations tended to reside in the central and lower parts of the county (which tend to have higher incomes), whereas those with lower-income destinations came mostly from northern CBGs (which tend to have lower incomes).
-Similar within-county dynamics happen in the other four counties.
+Similar within-county dynamics occur in the other four counties.
 Overall, this analysis illustrates the important demographic heterogeneity in migration which is concealed at the county level.
-Appendix F Privacy Protections
+Appendix D Case Study: New York City Housing Authority
+Ensuring that residents have access to affordable housing is an important and challenging policy goal [68].
+One way of achieving this is through the creation and management of public housing units: for example, the New York City Housing Authority (NYCHA) is the largest public housing authority in the US, providing affordable housing to over 500,000 authorized residents [69].
+We provide a case study illustrating how the MIGRATE dataset can be used to understand which residents make use of public housing units, thereby illuminating the effect of this component of affordable housing policy.
+Fig.
+S7: We use MIGRATE to understand migration to and from New York City Housing Authority (NYCHA) developments.
+(a) A map of the 114 Census Block Groups (marked in red) where all residential units are owned by NYCHA (i.e., where all the population lives in a NYCHA development).
+(b) Time series of in- and out-migration rates for NYCHA CBGs (red line), as well as for CBGs nearby (within 250 meters of a NYCHA development; green line) and non-NYCHA CBGs within New York City (gray line).
+NYCHA CBGs have consistently lower in- and out-migration rates, implying a less mobile population.
+(c) Demographics of previous residence CBG for in-movers into NYCHA buildings as well as for the other CBG groups considered, aggregated across the decade (error bars represent standard deviation across years).
+(d) Demographics of next residence CBG for out-movers.
+We identify all 114 New York City Census Block Groups whose residential units are entirely owned by NYCHA using New York City’s Primary Land Use Tax Lot Output (PLUTO) dataset [70] (Fig.
+S7a).
+We compare migration patterns for these CBGs, which we refer to as “NYCHA CBGs”, to migration patterns of two other groups of non-NYCHA CBGs: (i) nearby CBGs (within 250 meters of a NYCHA development) and (ii) all non-NYCHA CBGs in New York City.
+[^10] We excluded from the analysis all 345 CBGs (about 5.3% of all New York City CBGs) where some but not all of the residential units are owned by NCYHA.
+NYCHA CBGs have had consistently lower in-migration and out-migration rates than the two comparison groups of CBGs throughout the decade (Fig.
+S7b), implying a less mobile population.
+This is consistent with prior findings that residents of public housing and recipients of rental benefits have longer tenancies than average New York City residents [69, 71].
+To get a sense of the residents being served by NYCHA and their migration trajectories, we examine the demographics of the CBGs where NYCHA in-movers move from (Fig.
+S7c) and where NYCHA out-movers move to (Fig.
+S7d).
+Compared to non-NYCHA in-movers, NYCHA in-movers are likelier to move from plurality Black or Hispanic CBGs, as well as bottom income quartile CBGs.
+The same is true of NYCHA out-movers: they are also more likely to move to plurality Black or Hispanic CBGs, and bottom-income-quartile CBGs, than are non-NYCHA out-movers.
+The CBGs that NYCHA movers move to are higher-income than the CBGs they move from, providing evidence of upward mobility.
+NYCHA out-movers are also more upwardly mobile than non-NYCHA movers: 82% of NYCHA out-movers move to CBGs of higher median income, compared to 55% of all non-NYCHA out-movers.
+We also quantify whether NYCHA in-movers and out-movers move within NYCHA, within New York City, and within New York State.
+There is a significant amount of shuffling within the NYCHA system: 14% of in-movers to a NYCHA CBG came from another NYCHA CBG, and 13% of out-movers from a NYCHA building go to another NYCHA CBG.
+This shuffling may be due in part to mandated renovations or building closures that result in temporary resident relocations by NYCHA [72, 73, 74].
+On average across the decade that we study, NYCHA out-movers are likelier than non-NYCHA movers to remain within New York City (74% vs.
+65%); however, this gap has decreased in recent years.
+Overall, this analysis highlights how MIGRATE can provide insights into which residents are served by public housing policy.
+Beyond this case study of New York City, this approach could be extended to study public housing in other cities, or to study the effects of public housing foreclosures—which has been a topic of interest among other migration researchers [18].
+Appendix E Privacy Protections
 In this section we detail the measures we took to preserve the privacy of individuals in the MIGRATE data release.

### v3

@@
 Abstract
-Fine-grained migration data illuminate important demographic, environmental, and health phenomena.
-However, migration datasets within the United States remain lacking: publicly available Census data are neither spatially nor temporally granular, and proprietary data have higher resolution but demographic and other biases.
-To address these limitations, we develop a scalable iterative-proportional-fitting based method that reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data.
-We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 that captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data.
-These estimates are highly correlated with external ground-truth datasets, and improve accuracy and reduce bias relative to raw proprietary data.
-We use MIGRATE to analyze both national and local migration patterns.
-Nationally, we document temporal and demographic variation in homophily, upward mobility, and moving distance: for example, we find that people are increasingly likely to move to top-income-quartile CBGs and identify racial disparities in upward mobility.
-We also show that MIGRATE can illuminate important local migration patterns, including out-migration in response to California wildfires, that are invisible in coarser previous datasets.
-We publicly release MIGRATE to provide a resource for migration research in the social, environmental, and health sciences.
+Fine-grained migration data illuminate demographic, environmental, and health phenomena.
+However, United States migration data have serious drawbacks: public data lack spatiotemporal granularity, and higher-resolution proprietary data suffer from multiple biases.
+To address this, we develop a method that fuses high-resolution proprietary data with coarse Census data to create MIGRATE: annual migration matrices capturing flows between 47.4 billion US Census Block Group pairs—approximately four thousand times the spatial resolution of current public data.
+Our estimates are highly correlated with external ground-truth datasets and improve accuracy relative to raw proprietary data.
+We use MIGRATE to analyze national and local migration patterns.
+Nationally, we document demographic and temporal variation in homophily, upward mobility, and moving distance—for example, rising moves into top-income-quartile block groups and racial disparities in upward mobility.
+Locally, MIGRATE reveals patterns such as wildfire-driven out-migration that are invisible in coarser previous data.
+We release MIGRATE as a resource for migration researchers.
 1 Introduction
 Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena.
-Migration data illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; consequences of conflicts [8]; effects of the COVID-19 pandemic [9]; and political polarization [10].
+Migration data illuminate responses to environmental disasters and climate change hoffmann_drought_2024, kaysen_open_2025, mcconnell_rare_2024, alexander2019impact; responses to economic stresses and opportunities diamond_effect_2020, wilkerson_warmth_2010, niva_worlds_2023, boucherie_decoupling_2025; patterns of social change depillis_for_2024; consequences of conflicts meta-migration, leasure2023nowcasting; effects of the COVID-19 pandemic ramani_how_2024; housing instability phillips_measuring_2020, gershenson2024fracking; urban-suburban migration patterns golding2020tracking, reia2022spatial; and political polarization kaysen_millions_2024.
 But migration datasets within the United States have serious limitations.
-The most fine-grained freely available datasets track migration at the county level [11].
+The most fine-grained publicly available national datasets track migration at the county level county_to_county_data.
 While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level.
-For example, research on flood-risk-induced migration uses models at the much more granular Census block level [12], arguing that this spatial granularity is necessary due to the highly localized nature of flood risk [13].
-Research on neighborhoods that are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [14].
-Movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 [15].
-Publicly available, county-level migration datasets render these patterns entirely invisible.
+For example, research on flood-risk-induced migration uses models at the much more granular Census block level shu2023integrating, arguing that this spatial granularity is necessary due to the highly localized nature of flood risk columbia2024climatemigration.
+Other research on climate-induced vulnerability, eviction or displacement similarly models risk at a sub-county (Census tract) level Tedesco2021, gershenson2024fracking or even at the household level rising2022missing.
+Research on housing instability often models Census Block Groups freedman_living_2022 or even specific building complexes rutan_concentrated_2021.
+All of these applications testify to the need for highly granular data to facilitate accurate study of many important migration-related phenomena.
+Additionally, movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 kerns-damore_migration_2023.
+Publicly available, county-level migration datasets are too coarse to study these important migration patterns.
 Proprietary migration datasets offer greater granularity.
-Data aggregators like Infutor [16] combine many data sources — including voter files, property deeds, credit files, and phone books [17] — to attempt to infer address histories for individual-level movers.
-Such datasets have been widely used because they are extremely temporally and spatially fine-grained [18, 4, 19, 20, 21].
-Previous work also suggests these data do contain valuable signal which correlates with external datasets, as validated by cross-referencing the data with hurricanes or public housing foreclosures [18], and with marginal population counts in the region of interest [17].
-However, these proprietary datasets have three major disadvantages.
+Data aggregators like Infutor infutor combine many data sources — including voter files, property deeds, credit files, and phone books diamond_effects_2019, phillips2024homelessness, boar2023consumption — to attempt to infer address histories for individual-level movers.
+Such datasets have been widely used because they are extremely temporally and spatially fine-grained phillips_measuring_2020, diamond_effect_2020, bernstein2022contribution, asquith2019supply, qian2021effects, phillips2024homelessness.
+A long literature which provides more recent or granular migration estimates using non-traditional data sources, including social media and digital advertisement data, testifies to the power of new data sources meta-migration, alexander2019impact, alexander2022combining, leasure2023nowcasting, rampazzo2025assessing.
+Previous work also suggests that address history data do contain valuable signal which correlates with external datasets, as validated by cross-referencing the data with hurricanes or public housing foreclosures phillips_measuring_2020, and with marginal population counts in the region of interest diamond_effects_2019.
+However, proprietary address history datasets have three major disadvantages.
 First, they are not publicly available, limiting their utility to researchers.
@@
 Third, and most fundamentally, they combine multiple imperfect data sources using proprietary algorithms, and thus contain noise and biases.
-For example, Infutor and other consumer record datasets have been shown to overrepresent higher-income and majority-group populations in some settings [22].
-To address the limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates that combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data.
-To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm that repeatedly scales a noisy matrix to reconcile it with constraints [23, 24, 25], to reconcile the raw Infutor migration matrix with the more reliable Census constraints.
-We accommodate uncertainty in address mapping and incorporate data on births, deaths, and immigration.
-The output of our method is a set of yearly inferred migration matrices at the Census block group (CBG) level from 2010-2019 covering the 50 states and the District of Columbia.
-Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows publicly available on the 5-year level, and 18 million times more granular than the state-state flows publicly available on the 1-year level.
+For example, Infutor and other consumer record datasets have been shown to overrepresent higher-income and majority-group populations in some settings ramiller_residential_2024.
 Figure 1: MIGRATE estimates.
@@
 Out-migration rates for CBGs with fewer than 100 people are omitted.
-We comprehensively validate MIGRATE by comparing it to external data sources.
-We show that our estimates correlate well with fine-grained Census and ACS data on population counts and migration flows.
-We show that MIGRATE improves accuracy relative to raw Infutor data: in particular, it reduces the deviation from ground truth ACS county-to-county migration estimates by 90%.
-We document systematic demographic biases in the Infutor data, which overcounts rural, older, white, and home-owning populations, and show that MIGRATE almost entirely eliminates these biases.
+To address the limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates that combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data.
+To produce our estimates, we develop a data fusion methodimbens_combining_1994, guan2025datafusionhighresolutionestimation based on iterative proportional fitting (IPF)IPF, chang2024inferring, chang2021mobility to reconcile the raw migration data from data aggregator Infutor with the more reliable Census constraints.
+The output of our method is a set of yearly inferred United States migration matrices at the Census Block Group (CBG) level from 2010-2019.
+Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows publicly available on the 5-year level, and 18 million times more granular than the state-state flows publicly available on the 1-year level.
+We comprehensively validate MIGRATE by comparing to external data sources, showing that it correlates well with ground-truth population counts and migration flows, and improves accuracy and reduces demographic bias compared to raw Infutor data, which overcounts rural, older, white, and home-owning populations.
 We then use MIGRATE to analyze both national and local migration.
@@
 Locally, we show that MIGRATE can illuminate important migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data.
-Collectively, these analyses illustrate how MIGRATE reveals important migration patterns at both the national and local levels.
-To provide a foundation for more precise migration research in the social, environmental, and health sciences, we release MIGRATE for other researchers at migrate.tech.cornell.edu.
-2 Results
-2.1 Creating MIGRATE
+To provide a foundation for more precise migration research in the social, environmental, urban, and health sciences, we release MIGRATE for non-profit research use at migrate.tech.cornell.edu.
+2 Creating MIGRATE
 Here we provide a high-level summary of our method for inferring fine-grained migration matrices by harmonizing Census and Infutor data.
 In the Methods Section, we provide full details of processing the Infutor dataset (§M.1), processing Census data (§M.2), and harmonizing both data sources (§M.3).
-Preprocessing the raw Infutor dataset
+2.1 Preprocessing the raw Infutor dataset
 We first preprocess the raw Infutor data into yearly migration matrices.
-The raw Infutor data consists of sequences of addresses for individual people.
-We use this information to compute the number of people moving between each pair of geographic areas.
-In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior [^1] We correct populations to account for birth, deaths, and international migration.
-Full details in the Methods Section (§M.2)..
-We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Table M1.
+The raw Infutor data consists of sequences of addresses for individual people [^1] For a subset of individuals, Infutor also provides estimated demographic information (e.g., age, gender, and socioeconomic information).
+We do not incorporate this information in our analysis for two reasons: first, complete demographic data is only available for about half of all individuals; second, the reliability of estimated demographic information is concerning—as the data lacks documentation, we should not discard the possibility that demographic information has been imputed based on addresses—and contains known and potential biases ramiller_residential_2024, 10.1093/pnasnexus/pgaf027.
+Analogously, we do not use Census information on demographic subgroups when harmonizing with Census data; we only make use of population counts and flows..
+We use these address sequences to compute the number of people moving between each pair of geographic areas.
+In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior.
+We provide a conceptual overview of the preprocessing pipeline here and full details in the Methods Section M.1.
+We first construct cleaned monthly address histories for each individual in the Infutor dataset by reconciling inconsistencies in address start/end dates, discarding unreliable records (e.g., postal boxes), and modeling uncertainty when multiple addresses are active.
+These monthly address distributions are then aggregated in a way that simulates ACS survey responses about residence one year prior, producing annual migration flow estimates between pairs of addresses.
+Finally, to convert the address-address migration matrices to CBG-CBG migration matrices, each address is mapped to one or more CBGs.
+When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder R-arcgeocoder, us_census_bureau_census_2024.
+We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results.
+The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code.
+We are able to map 99.21% of all addresses in Infutor to CBGs.
+The final output of this preprocessing procedure is the Infutor yearly migration matrices $E^{(t)}$, which capture migration counts between pairs of CBGs as estimated from Infutor data alone.
+We provide summary statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Table 2.
 Individuals with active records (any time in 2010-2019) 374,217,253
@@
 Address records per active individual (median) 2
-Table 1: Summary statistics for raw Infutor data and Census data during the period of interest (2010-2019).
+Table 1: Summary statistics for raw Infutor data and Census data (2010-2019).
 We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period.
 Comparing the average yearly number of active individuals in the Infutor data to the Census population (second row) shows that Infutor under-counts the population.
-(Table M1 provides the number of active individuals in each year from 2010-2019).
+(Table 2 provides the number of active individuals in each year from 2010-2019).
 Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
-Preprocessing Infutor requires mapping addresses to geographic areas.
-When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder [26, 27].
-We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results.
-The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code.
-We are able to map 99.21% of all addresses in Infutor to CBGs.
-We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in the Methods Section (§M.1).
-Harmonizing the Infutor and Census datasets
+2.2 Harmonizing the Infutor and Census datasets
 Having produced the raw Infutor yearly migration matrices $E^{(t)}$, we next reconcile them with more reliable but less granular Census data by rescaling selected entries of $E^{(t)}$.
-For example, to reconcile $E^{(t)}$ with Census county populations at time $t$, we multiply the rows corresponding to each county by a scaling factor to match the Census population in that county.
 We rescale to different Census datasets in sequence, rather than simultaneously, because the datasets are not totally consistent with each other.
-Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations; then the 1-year counts of movers and non-movers by state; then the 1-year state-to-state flows; and finally the 1-year county populations.
-The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them: while CBG populations and state flows are estimated by the American Community Survey (ACS) and often contain large sampling errors, county populations are estimated by the Population Estimates Program (PEP) using more precise large-scale administrative datasets.
-We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (Appendix A.2).
-To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [23, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges.
+Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations from the Census and 5-year American Community Survey (ACS); then the 1-year ACS counts of movers and non-movers by state; then the 1-year ACS state-to-state flows; and finally the 1-year county populations from the Population Estimates Program (PEP).
+The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them.
+While CBG populations and state flows are estimated by the ACS and often contain large sampling errors (see our discussion in Methods Section M.2), county populations are estimated by PEP using more precise large-scale administrative datasets.
+We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (see Supplementary Table 1).
+To perform the final rescaling (matching 1-year county populations) we apply an iterative procedure based on the classical IPF algorithm IPF, chang2024inferring: specifically, we scale blocks of rows and columns to agree with the annual county populations in years t1t-1 and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges.
+Only the final rescaling is performed using full IPF; the other rescalings, to CBG populations, 1-year counts of movers and non-movers, and 1-year state-to-state flows, are performed only once at the beginning of the process, as opposed to iteratively, to avoid overfitting to these more noisily estimated datasets.
 The resulting migration matrices constitute our MIGRATE estimates.
-The Methods Section (§M.3) provides additional methodological details on these scalings.
+In Methods Section M.3, we provide additional details on these scalings.
 Naive implementations of our harmonization algorithm impose prohibitive computation times due to the size of the matrices involved.
@@
 This granularity often reveals important patterns which are invisible at the county-county level, as we illustrate by zooming in on New York City and comparing the county-level rates obtained from ACS flows (Figure 1b) to the CBG-level out-migration rates inferred by MIGRATE (Figure 1c).
-2.2 Validating MIGRATE
+3 Validating MIGRATE
 We conduct extensive validations of the MIGRATE estimates.
@@
 We then investigate demographic biases in the raw Infutor estimates and show that MIGRATE reduces them (Figure 3).
+In Appendix A.2 we further verify, with experiments on semi-synthetic data, that our method recovers ground-truth flows under certain conditions of bias and noise.
 Figure 2: Validating the MIGRATE estimates.
-(a - c): MIGRATE estimates (y-axis) are highly correlated with Census data (x-axis), including (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population).
+(a - c): MIGRATE estimates (y-axis) are highly correlated with Census data (x-axis), including (a) Census populations at the tract, and block group (CBG) level, (b) movers between each pair of states and each pair of counties (excluding people who remain within the same state or county), and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population).
 For in-migration rates, correlations are weighted by state or county population, and points are sized by population.
-(d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data.
-We compute RMSE between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates.
-Bars show the mean reduction in RMSE across all years; error bars plot standard deviation across years.
+(d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data for population counts, movers between states and counties, and in-migration rate, respectively.
+We compute root mean squared error (RMSE) between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates.
+Bars show the mean reduction in RMSE across all data release years ($n=5$ for population and county-level migration datasets, $n=9$ for state-level migration datasets); error bars plot standard deviation across years.
 For the in-migration rates, RMSE is weighted by area population.
-MIGRATE estimates are highly correlated with ground-truth Census data
-MIGRATE estimates are well-correlated with all available Census measures (Figure 2a-b).
-We first assess correlations with Census datasets incorporated in our harmonization procedure, to confirm it indeed helps reconcile MIGRATE with Census data.
-We would expect these correlations to be high but not uniformly perfect, because the Census datasets are not totally consistent with each other.
-By design, MIGRATE estimates perfectly match state and county populations (because the last step in our procedure is to match to county populations).
-MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations [^2] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data.
-We only use ACS data products whose period completely overlaps with the 2010-2019 MIGRATE range..
-MIGRATE estimates are also highly correlated with state-to-state flows ($\rho=1$); this high correlation is driven in part by variation in state populations, as well as by the fact that most people do not move.
-Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix A.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
-Importantly, we further verify that our estimation procedure yields highly correlated estimates even with datasets which are not used in our estimation procedure.
-This constitutes a validation on held-out data, a standard check for statistical estimation methods [28].
-Specifically, we compare our MIGRATE estimates to ACS 5-year county-to-county flows, which are not incorporated in our estimation, and find they are highly correlated with our estimates ($\rho=1$).
-This high correlation is driven in part by the fact that most people do not move; hence, in Appendix A.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$).
-Another contributor to the high correlation in county-county flows is variation in county size, so we separately compute correlations with county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) which normalize for county size.
-Correlations remain high (population-weighted $\rho=0.715$).
-As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation: for example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census block group populations).
-See Appendix A.2 for full results.
-Overall, these validations demonstrate that MIGRATE estimates are highly correlated with external ground-truth Census data.
-This is not merely because of variation in state or county size (since it remains true when examining in-migration rates, which normalize for state or county size, as well as when looking at the tract and CBG level).
-Nor is it due merely to the fact that most people do not move (since it remains true when examining correlations for movers specifically).
-Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets that are not used in our estimation procedure.
-MIGRATE estimates reduce error relative to raw Infutor data
-We show that MIGRATE estimates also increase agreement with Census datasets compared to using the Infutor data.
-We note that Infutor systematically undercounts the population (Table 1); to compare to Infutor as generously as possible, and in particular to ensure that MIGRATE does not reduce error due to trivial rescalings, we multiply each raw Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population prior to conducting these comparisons.
-Figures 2c and 2d report the reduction in error that MIGRATE estimates achieve relative to Infutor data.
+3.1 MIGRATE estimates are highly correlated with ground-truth Census data
+MIGRATE estimates are well-correlated with all available Census measures of population counts and flows (Figure 2a-c).
+By design, MIGRATE estimates perfectly match state and county populations (because the last step in our harmonization procedure is to match to county populations).
+MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations (Figure 2a) [^2] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data, using only ACS data products whose time period completely overlaps with the 2010-2019 MIGRATE range..
+MIGRATE estimates of the number of movers between each pair of states and each pair of counties are also highly correlated with ACS estimates ($\rho=0.998$ and $0.957$ for states and counties respectively; Figure 2b) [^3] We exclude people who remain within the same state or county from this calculation, because most people do not move, which artificially inflates the correlation.
+Supplementary Table 3(b) reports the correlation without this exclusion, which is nearly perfect..
+Finally, MIGRATE estimates of state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) are well-correlated with ACS estimates (Figure 2c; $\rho=0.987$ and 0.715 for state and county in-migration rates, respectively, weighting by area population).
+Supplementary Table 1(a) reports correlations with additional Census quantities (e.g., in-migration counts as opposed to rates) which remain high.
+Overall, these validations demonstrate that MIGRATE estimates are highly correlated with ground-truth Census data.
+This is not merely because of variation in state or county populations, since it remains true when examining in-migration rates, which normalize for population, as well as when looking at tracts and CBGs, which do not vary as significantly in population.
+Nor is it due merely to the fact that most people do not move, since we examine correlations for movers specifically, as well as correlations with in-migration rates.
+Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets that are not used in our harmonization procedure, a standard check for statistical estimation methods hastie_elements_2009444Even for some Census datasets that are used in our harmonization procedure, we would expect correlations with MIGRATE to be high but not perfect, because the Census datasets are not totally consistent with each other..
+Specifically, MIGRATE estimates remain highly correlated with county-county mover counts and county in-migration rates, which are not used in our estimation procedure.
+As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation.
+For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with data below county level ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census Block Group populations).
+See Supplementary Table 1(a) for full results.
+3.2 MIGRATE estimates reduce error relative to raw Infutor data
+We show that MIGRATE estimates also increase agreement with Census datasets compared to using the Infutor data alone.
+We note that Infutor systematically undercounts the population, as shown in Table 1; to compare to Infutor as generously as possible, and in particular to ensure that MIGRATE does not reduce error due to trivial rescalings, we multiply each raw Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population prior to conducting these comparisons.
+Figures 2d-f report the reduction in error that MIGRATE estimates achieve relative to Infutor data.
 We compute root mean squared error (RMSE) between (1) MIGRATE estimates and ground-truth Census data and (2) Infutor data and ground-truth Census data, and report the reduction in RMSE from using MIGRATE estimates.
-We show average reduction in RMSE across all data releases between 2010 and 2019; error bars represent standard deviation across these releases.
+We report average reduction in RMSE across all data releases between 2010 and 2019; error bars represent standard deviation across these releases.
 MIGRATE estimates eliminate error in state and county populations because they are constructed to match county-level estimates.
-Relative to raw Infutor data, MIGRATE estimates reduce error in Census tract populations by an average of $85.9\%$ and error in CBG populations by $83.6\%$.
-MIGRATE estimates reduce error in state-to-state flows by $89.5\%$ and error in state in-migration rates by $83.3\%$.
-We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix A.1.
-We again confirm that we still see these reductions in error on held-out data which is not used in our estimation procedure.
-Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$; neither of these datasets was used in producing MIGRATE estimates.
-We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix A.2).
+They also reduce error in Census Tract and CBG populations by an average of $85.9\%$ and $83.6\%$ respectively (Figure 2d); in state-to-state movers and county-to-county movers by $87.5\%$ and $42.3\%$ (Figure 2e); and in state in-migration and county in-migration rates by $87.3\%$ and $51.8\%$ (Figure 2f).
+These gains persist even when using held-out datasets that are not used in estimating MIGRATE (namely, county-to-county flows and county in-migration rates).
+We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Supplementary Table 1(b)).
 Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data.
-MIGRATE estimates reduce demographic biases present in raw Infutor data
+We provide error reductions for additional metrics (all flows, non-movers, and in-migration counts as opposed to rates) in Supplementary Table 1(b), and conduct additional validations on synthetic data (Appendix A.2).
+3.3 MIGRATE estimates reduce demographic biases present in raw Infutor data
 Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates.
@@
 (c) Comparison of demographic bias in Infutor (purple) and MIGRATE (green).
-Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data.
+Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data, averaged over $n=5$ population data releases (American Community Survey 5-year estimates, 2015 through 2018) Error bars represent standard deviations across these releases.
 MIGRATE greatly reduces biases for all demographic subgroups.
 The Infutor data displays geographic and demographic biases (Figure 3).
-Specifically, it overrepresents populations in the counties in the northeastern United States while underrepresenting populations in the southwest (Figure 3a).
-[^3] For this analysis, as above, we multiply each Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population to avoid deviations due to trivial rescalings.
+Specifically, it overrepresents populations in the counties in the northeastern United States while underrepresenting populations in the southwest (Figure 3a) [^5] For this analysis, as above, we multiply each Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population to avoid deviations due to trivial rescalings..
 Errors plotted are averages of county population errors across all 10 years of data from 2010-2019.
@@
 Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups.
-These biases are consistent with biases found in previous research [22], as we discuss in detail in Appendix A.3, and could propagate into biases in downstream analyses of inequality and other topics.
+These biases are consistent with biases found in previous research ramiller_residential_2024, as we discuss in detail in Appendix A.3, and could propagate into biases in downstream analyses of inequality and other topics.
 MIGRATE estimates reduce biases in Infutor data.
@@
 The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases.
-We present analyses for additional demographic groups, finding similar results, in Appendix A.4.
-2.3 Analysis of national migration patterns
+We present analyses for additional demographic groups, including immigrant populations, in Appendix A.4.
+4 Analysis of national migration patterns
 Figure 4: National migration statistics.
-(a) Flows between ten types of CBGs — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile.
+MIGRATE can be used to reveal fine-grained patterns of mobility across the United States.
+(a) Flows between ten types of Census Block Groups (CBGs) — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile.
 Rows correspond to the origin CBG, and columns to the destination CBG; for example, the top left entry indicates that 90% of movers from plurality white CBGs move to plurality white CBGs.
@@
 We use MIGRATE to analyze national patterns of migration from 2010-2019.
-The spatial granularity of the data affords a unique opportunity to study demographic variation in migration patterns — for example, differences in migration between higher-income and lower-income CBGs — that may be obscured by county or state-level data, which show far less demographic variation.
-We hence divide CBGs into 10 (overlapping) categories — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile — and stratify the migration statistics we compute by these ten categories [^4] To determine the category of each CBG every year, we use the most recent ACS 5-year demographic estimates.
-For example, to classify CBGs into plurality races to study moves in the 2010-2011 period, we use the ACS 2006-2010 race and ethnicity estimates..
-Figure 4a plots flows between the ten categories — for example, the proportion of movers from top-income-quartile CBGs who move to top-income-quartile CBGs.
-This reveals substantial homophily in migration: out-movers from all ten CBG categories are more likely than movers as a whole to move to CBGs of the same category.
-However, the strength of this homophily varies across categories.
-Movers from plurality Black, Asian, and Hispanic CBGs exhibit strong homophily: they are 5.5$\times$, 14.6$\times$, and 4.3$\times$ likelier than movers as a whole to move to CBGs with the same plurality race group (Figure S3a reports relative rates for all groups).
-We also observe income homophily, particularly for movers within top and bottom quartile CBGs: movers from bottom-income-quartile CBGs are nearly twice as likely as movers as a whole to move to CBGs in the bottom income quartile (34% versus 18% for movers as a whole); movers from top-income-quartile CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs (53% versus 32%).
-In Figure S3b, we confirm that this homophily does not occur merely because many moves are local and demographics are spatially correlated: we also observe homophily when restricting to long-distance (out-of-county) moves.
-Figure 4a also shows that people are likelier to move to CBGs in higher income quartiles: 32% of moves are to top-income-quartile CBGs, while only 18% are to bottom-income-quartile CBGs.
-This trend becomes more pronounced over the decade we study (Figure S4d) and is only partially explained by the fact that top-income-quartile CBGs account for a larger share of the population (29%) than bottom-income-quartile CBGs (21%).
-There are also racial disparities: movers from plurality Asian CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs; movers from plurality Black CBGs are 2.0$\times$ likelier than movers as a whole to move to bottom-income-quartile CBGs.
+The spatial granularity of the data affords an opportunity to study demographic variation in migration patterns — for example, differences in migration between higher-income and lower-income CBGs — that may be obscured by county or state-level data, which show far less demographic variation.
+We hence divide CBGs into 10 (overlapping) categories — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile — and stratify the migration statistics we compute by these ten categories.
+To determine the category of each CBG every year, we use the most recent ACS 5-year demographic estimates.
+For example, to classify CBGs by their plurality race group to study moves in the 2010-2011 period, we use the ACS 2006-2010 race and ethnicity estimates.
+[…]
+[…]
 Figure 4b investigates whether these racial disparities persist when controlling for origin CBG median income: in particular, plotting the share of movers moving to a higher-income CBG when controlling for the mover’s origin CBG median income decile.
@@
 For example, movers from fifth-income-decile, plurality Asian CBGs have a 71% chance of moving to a higher-income CBG; movers from fifth-income-decile, plurality Black CBGs have a 53% chance.
-Figure S4 shows that the same racial disparities occur when controlling for CBG median income percentile, as opposed to decile; or when examining the probability of moving to a top- or bottom-income-quartile CBG, as opposed to moving to a higher-income CBG.
+Supplementary Figure 5 shows that the same racial disparities occur when controlling for CBG median income percentile, as opposed to decile; or when examining the probability of moving to a top- or bottom-income-quartile CBG, as opposed to moving to a higher-income CBG.
 Overall, we find robust and substantial racial disparities in income of destination CBG that persist even conditional on income of origin CBG.
@@
 Figure 4c also highlights demographic variation in these statistics, revealing that movers from plurality white CBGs, rural CBGs, and higher-income CBGs are likelier to move long distances (more than 50 miles).
-Appendix B.4 reports additional migration distance statistics stratified by geographic boundary (i.e., whether movers remain within the same tract, county, or state) and shows that migration distance increases over the decade we study.
+In Appendix B.4, we report additional migration distance statistics stratified by geographic boundary (i.e., whether movers remain within the same tract, county, or state) and show that migration distance increases over the decade we study.
 Overall, these results demonstrate that fine-grained migration data illuminates important demographic variation in homophily, upward mobility, and migration distance.
-2.4 Analysis of local migration patterns
+Future research could stratify migration flows by additional characteristics available in Census data: for example, one might study the migration patterns of immigrants by analyzing flows from areas with larger proportions of immigrants, or larger proportions of residents with a given country of birth—both of these are datasets available at the sub-county level via ACS.
+However, we note that such analyses require significant heterogeneity across CBGs (or other Census areas) in the demographic trait being studied; for example, it would not be possible to reliably disaggregate migration patterns by gender using this method, since gender proportions remain relatively stable across Census areas.
+5 Analysis of local migration patterns
+In addition to using MIGRATE to analyze national migration trends, we use it to study local migration patterns: specifically, migration in response to wildfires in California.
+Natural disasters, including wildfires, are known drivers of human mobility mcconnell_rare_2024, mcconnell2021effects.
+There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area.
+In many US states, including California, wildfire risk increased from 2010-2019 modaresi_rad_human_2023, and is expected to further increase due to climate change Barbero_2015.
+Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters din_administrative-data.
 Figure 5: Migration in response to California wildfires.
 (a) Out-migration following the Camp fire (2018; top) and Tubbs fire (2017; bottom).
-Red boundaries plot fire perimeters; black lines plot county boundaries; CBGs are colored by domestic out-migration rate in MIGRATE.
+Red boundaries plot fire perimeters; black lines plot county boundaries; Census Block Groups (CBGs) are colored by domestic out-migration rate in MIGRATE.
 Out-migration rates exceed 50% in many CBGs within the fire perimeters.
@@
 Out-migration rates in the year after the fire are higher in CBGs within the fire perimeter (red line) than in groups of CBGs outside the fire perimeter (other lines), including those neighboring the fire perimeter, those in affected counties, or those within California.
-(c) Out-migration rates in ACS 5-year county-to-county data remain relatively constant over time.
-In addition to using MIGRATE to analyze national migration trends, we use it to study local migration patterns: specifically, migration in response to wildfires in California.
-Natural disasters, including wildfires, are known drivers of human mobility [3, 29].
-There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area.
-In many US states, including California, wildfire risk increased from 2010-2019 [30], and is expected to further increase due to climate change [31].
-Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters [32].
+(c) Out-migration rates in the American Community Survey (ACS) 5-year county-to-county data remain relatively constant over time.
 Post-wildfire migration estimates can inform disaster response and long-term planning in multiple ways.
 First, migration data is crucial for policymaking in the aftermath of a wildfire.
-It helps guide the allocation of housing, public health, and other resources to support displaced residents while minimizing strain on housing markets in receiving areas [33, 29, 34, 35].
-Data on out-migration rates from affected areas can also inform decisions about whether and how to rebuild in fire-prone regions [36, 37].
+It helps guide the allocation of housing, public health, and other resources to support displaced residents while minimizing strain on housing markets in receiving areas rosenthal2021health, mcconnell2021effects, hopfer2024repeat, isaac2023differences.
+Data on out-migration rates from affected areas can also inform decisions about whether and how to rebuild in fire-prone regions mcwethy2019rethinking, mach2021reframing.
 Second, high-resolution migration data is important for future wildfire planning.
-As populations relocate in response to shifting wildfire risks, migration patterns can refine policymakers’ estimates of future risk and guide the regulation of housing and insurance markets [38, 39, 40, 41, 42].
-While some households may move to reduce their exposure, others may become “trapped” in high-risk areas due to financial, social, or logistical barriers [43, 44, 45].
-Identifying these communities through high-resolution migration data can help ensure that government support reaches those who need it most [46, 37].
-We use high-resolution fire perimeter data from the California Department of Forestry and Fire Protection [47].
+As populations relocate in response to shifting wildfire risks, migration patterns can refine policymakers’ estimates of future risk and guide the regulation of housing and insurance markets boomhower_how_2024, baylis_adapatation_2024, kramer_post-wildfire_2021, alexandre_rebuilding_2014, lee_post-disaster_2024.
+While some households may move to reduce their exposure, others may become “trapped” in high-risk areas due to financial, social, or logistical barriers bohra2014nonlinear, cattaneo2016migration, benveniste2024global.
+Identifying these communities through high-resolution migration data can help ensure that government support reaches those who need it most hino2017managed, mach2021reframing.
+We use high-resolution fire perimeter data from the California Department of Forestry and Fire Protection california_department_of_forestry_and_fire_protection_historical_2024.
 Fire perimeters are typically much smaller than county boundaries, suggesting that analyzing fire impacts may require the granularity of the MIGRATE estimates as opposed to the relatively coarse county-to-county flows.
 We analyze the two most destructive fires in California from 2010-2019 (as quantified by the number of structures lost to the fire): the Tubbs Fire in October 2017, and the Camp Fire in November 2018 (Figure 5).
-The Tubbs Fire occurred in Napa County and Sonoma County and destroyed at least 5,636 residential or commercial structures [48].
-The Camp Fire, about a year later, destroyed over 18,804 structures and damaged nearly 1,000 more in the Northern California county of Butte [49] [^5] The Camp Fire remains the most destructive wildfire in California as of early 2025; the Tubbs Fire has only been surpassed by the January 2025 Eaton and Palisades fires [50]..
-MIGRATE estimates reveal dramatic levels of out-migration for CBGs which lie within the fire perimeters (Figure 5a) in the year following the fires, often exceeding 50%.
-[^6] The fact that the out-migration rate is not 100% is likely due to a number of factors, including the often-preferred option of rebuilding and returning [40, 41].
-Looking at administrative address data from mail delivery after Camp Fire, for example, researchers found a much more significant uptick in short-term vacancy than in long-term vacancy [32].
-In contrast, CBGs that do not lie within the fire perimeters experience much lower out-migration rates.
-We systematically quantify these differences in Figure 5b, which compares CBGs within the fire perimeter to three sets of less-affected CBGs outside the perimeter: other CBGs neighboring the perimeter; other CBGs within the affected counties; and others within California as a whole.
-The out-migration rate in the year following the Camp Fire for CBGs within the fire perimeter is 46%, at least 3.1$\times$ that of less-affected CBG groups; the out-migration rate in the year following the Tubbs Fire for CBGs within the perimeter is 37%, at least 2.8$\times$ that of less-affected CBG groups.
-Our estimates of out-migration following the Camp Fire are similar to those in prior work [29].
+The Tubbs Fire occurred in Napa County and Sonoma County and destroyed at least 5,636 residential or commercial structures california_department_of_forestry_and_fire_protection_tubbs_2025.
+The Camp Fire, about a year later, destroyed over 18,804 structures and damaged nearly 1,000 more in the Northern California county of Butte california_department_of_forestry_and_fire_protection_camp_2025.
+The Camp Fire remains the most destructive wildfire in California as of early 2025; the Tubbs Fire has only been surpassed by the January 2025 Eaton and Palisades fires california_department_of_forestry_and_fire_protection_top_2025.
+[…]
 In contrast, publicly available county-level migration data obscure these dramatic out-migrations; Figure 5c shows that rates of out-migration in affected counties remain essentially flat.
@@
 (For this analysis, we compare to the relative change in the ACS 5-year dataset released the year following the fire from the ACS 5-year dataset the year before: for example, for the 2018 Camp Fire, we compare the 2015-2019 ACS estimates to the 2014-2018 ACS estimates.) The ACS 5-year population estimates, unlike MIGRATE estimates, also provide no insight into the destinations of out-movers.
-In Appendices C and D, we provide two additional analyses of local migration patterns that are impossible using only county-level data: we analyze socioeconomic variation in New York City out-mover destinations, and migration patterns for residents who live in public housing provided by the New York City Housing Authority as part of the city’s affordable housing policy.
+In Appendix C, we provide two additional analyses of local migration patterns that are impossible using county-level data: we analyze socioeconomic variation in New York City out-mover destinations, and migration patterns for residents who live in public housing provided by the New York City Housing Authority as part of the city’s affordable housing policy.
 Collectively, these analyses showcase how MIGRATE estimates reveal important, policy-relevant local migration patterns that traditional data sources conceal.
@@
 MIGRATE estimates correlate highly with external Census ground-truth datasets and reduce error and demographic bias relative to raw proprietary data.
-We discuss the measures taken to protect the privacy of all individuals in the dataset in Appendix E.
-Our estimates are available at migrate.tech.cornell.edu.
-We use MIGRATE to analyze both local and national patterns of migration.
-We show that MIGRATE can reveal important local migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data.
-Nationally, we analyze demographic and temporal variation in migration patterns.
-We find that people tend to move to CBGs of the same type across dimensions of race, income, and rural/urban status — for example, movers from plurality Black, Asian, and Hispanic CBGs are 15.5$\times$, 14.6$\times$, and 4.3$\times$ likelier to move to CBGs with the same plurality race group, consistent with prior work documenting the persistence of racial segregation [51, 52, 53].
-We document demographic and temporal variation in moving distance: movers from plurality white, rural, and higher-income CBGs are likelier to travel long distances, and moving distance increases over the decade we study, consistent with prior work [15].
+Our estimates are available for non-profit research use at migrate.tech.cornell.edu.
+We discuss the measures taken to protect the privacy of all individuals in the dataset in the Supplementary Information.
+[…]
 An important finding from our national analysis is racial and temporal variation in upward mobility.
@@
 We find large racial disparities in the likelihood of moving to higher-income CBGs, or top-income-quartile CBGs, even when controlling for income of origin CBG: movers from plurality Asian CBGs are likelier to move to a higher-income CBG, and movers from plurality Black CBGs are less likely.
-These findings are consistent with prior work documenting racial disparities in neighborhood attainment [54, 55, 56, 57] and generalize these findings to larger and more recent samples and additional racial groups.
+These findings are consistent with prior work documenting racial disparities in neighborhood attainment south2011metropolitan, pais2012metropolitan, south2016neighborhood, quillian2015comparison and generalize these findings to larger and more recent samples and additional racial groups.
 Overall, we document demographic and temporal variation in national migration trends using a recent, large-scale, and highly granular dataset.
@@
 First, while we use multiple ground-truth Census data sources to reduce the biases we document in the raw Infutor data, uncorrected biases likely remain.
-For example, while we correct biases in CBG populations, we cannot correct sub-CBG-level population biases, or biases in CBG-CBG flows.
+For example, while we correct biases in CBG populations, we cannot correct sub-CBG-level population biases, or biases in CBG-CBG flows; practitioners should thus not assume data at this level contain no residual biases.
+Data at very fine-grained levels will also be noisier than data at more aggregated levels; while we provide practitioners with fine-grained data to maximize flexibility, we recommend aggregating up to less fine-grained levels if granularity is not required for analysis.
 Second, while we show that MIGRATE estimates are highly correlated with external Census datasets (including held-out datasets that are not used to produce our estimates), these validations have imperfections.
-As we further describe in the Methods Section (§M.2), Census datasets themselves have biases; can possess significant margins of error, particularly at fine spatial scales; and are not perfectly consistent with each other.
-Another challenge in validation is the lack of a ground-truth CBG-CBG migration matrix against which to validate; publicly available Census datasets do not afford this level of granularity, and non-Census data sources (e.g., voter files) have significant biases and measure substantively different populations than the one we seek to model [58, 59].
-Overall, MIGRATE should be viewed as a new data source which like all data possesses limitations but which nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
-We hope these improvements will facilitate a wide range of further migration-related analyses.
-As our analyses illustrate, MIGRATE reveals patterns which remain invisible in publicly available, county-level migration data.
-We hope MIGRATE will facilitate more precise study of many migration-related phenomena across the social, health, and environmental sciences.
+As we further describe in the Methods Section, Census datasets themselves have biases; are not perfectly consistent with each other; and can possess significant margins of error, particularly at fine spatial scales.
+To mitigate these concerns, we conduct our harmonization process using datasets with relatively low margins of error and also prioritize fitting to Census datasets that are more precisely estimated (both detailed in the Methods Section).
+Another challenge in validation is the lack of a ground-truth CBG-CBG migration matrix against which to validate; publicly available Census datasets do not afford this level of granularity, and non-Census data sources (e.g., voter files) have significant biases and measure substantively different populations than the one we seek to model fraga2018using, kim_when_2022, 10.1093/pnasnexus/pgaf027.
+Overall, MIGRATE should be viewed as a migration data source that like all data possesses limitations but that nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
+We hope these improvements will enable a wide range of further migration-related analyses.
+As our analyses illustrate, MIGRATE reveals patterns which cannot be observed in publicly available, county-level migration data.
+We release MIGRATE as a resource to facilitate more precise study of many migration-related phenomena across the social, health, urban, and environmental sciences.
 Methods
@@
 Infutor provides data in tabular form.
-Each row in the data provides data for one individual, which includes a list of known addresses, along with the date when the individual is first observed at the address (which we refer to the effective date below), and the first and last date that the individual is observed in the data (which we refer to as the listed start date and the listed end date, respectively).
-[^7] For a subset of individuals, Infutor also provides estimated demographic information (e.g., age, gender, and socioeconomic information).
-We do not incorporate this information in our analysis for two reasons: first, complete demographic data is only available for about half of all individuals; second, the reliability of estimated demographic information is concerning—as the data lacks documentation, we should not discard the possibility that demographic information has been imputed based on addresses—and contains known biases [22].
+Each row in the data provides data for one individual, which includes a list of known addresses, along with the date when the individual is first observed at the address (which we refer to as the effective date below), and the first and last date that the individual is observed in the data (which we refer to as the listed start date and thelisted end date, respectively).
 Dates are listed as month and year.
@@
 We describe the steps taken to (1) identify and clean Infutor data records within the scope of MIGRATE, (2) process the Infutor data into matrices documenting yearly flows between address pairs, and (3) map these address-level matrices to the CBG-level matrices $E^{(t)}$ mentioned in the main text.
-Table M1 summarizes the raw and processed data.
+Table 2 summarizes the raw and processed data.
 Cleaning address histories
@@
 We similarly define their reconciled end date as the maximum of their last effective date, and their listed end date.
-(For example, if the effective date list for an individual was [January 2013, January 2016], their listed start date was January 2014, and their listed end date was January 2017, their reconciled start date would be January 2013, and their reconciled end date would be January 2017.) Finally, we define their interval of activity as [reconciled start date -- 1 year, reconciled end date $+$ 1 year].
+(For example, if the effective date list for an individual was [January 2013, January 2016], their listed start date was January 2014, and their listed end date was January 2017, their reconciled start date would be January 2013, and their reconciled end date would be January 2017.) Finally, we define their interval of activity as [reconciled start date - 1 year, reconciled end date ++ 1 year].
 We use 1-year padding to reflect the annual granularity of the final dataset we produce and avoid discarding data in each yearly estimation process.
@@
 Finally, we create an yearly set of matrices $A^{(t)}$ with dimension $N_{\text{ADDRESSES}}\times N_{\text{ADDRESSES}}$ (i.e.
-a square matrix on the number of unique Infutor addresses) where each entry $A^{(t)}_{ij}$ corresponds to the expected number of individuals moving between addresses $i$ and $j$ from year $t-1$ to year $t$.
+a square matrix on the number of unique Infutor addresses) where each entry $A^{(t)}_{ij}$ corresponds to the expected number of individuals moving between addresses $i$ and $j$ from year t1t-1 to year $t$.
 These matrices will be further aggregated based on the location of each address.
 Processing address-address migration matrices into CBG-CBG migration matrices
-Fig.
-M1: Flowchart detailing the process of mapping addresses to CBGs.
+We transform the address-level migration matrices $A^{(t)}$ into CBG-level matrices $E^{(t)}$ by mapping each address either to a precise CBG where possible, or to a distribution over CBGs when only a ZIP code is available.
+We use the 2010 Census block group boundaries for all addresses to maintain geographic consistency.
+Figure 6 details this process.
+Figure 6: Geocoding overview.
+Flowchart detailing the process of mapping addresses to Census Block Groups (CBGs).
 We start with all Infutor addresses and classify them into five address types.
@@
 In total, we are able to map 99.21% of all addresses in the original Infutor data: 83.33% to a precise latitude and longitude, and 15.89% to a ZIP code.
-We transform the address-level migration matrices $A^{(t)}$ into CBG-level matrices $E^{(t)}$ by mapping each address either to a precise CBG where possible, or to a distribution over CBGs when only a ZIP code is available.
-We use the 2010 Census block group boundaries for all addresses to maintain geographic consistency.
-Figure M1 details this process.
 Our address mapping pipeline is divided in two parts: mapping addresses via geocoding, and mapping addresses with a ZIP code-to-CBG crosswalk.
-We initially attempt to use one of two state-of-the art geocoders to map addresses to latitudes and longitudes: the publicly-available Census Bureau geocoder [27] and ESRI’s ArcGIS geocoder [26].
+We initially attempt to use one of two state-of-the art geocoders to map addresses to latitudes and longitudes: the publicly-available Census Bureau geocoder us_census_bureau_census_2024 and ESRI’s ArcGIS geocoder R-arcgeocoder.
 We first attempt to geocode an address via the Census geocoder; if this algorithm does not find a match, we submit the address to ArcGIS.
@@
 Addresses representing rural routes and postal boxes are directly mapped via ZIP code, as well as incomplete addresses.
-We use the HUD ZIP-to-tract crosswalk, with the date closest available to the last seen date of an address to account for the fact that ZIP code definitions may vary and are not nested within Census geographies [60].
+We use the HUD ZIP-to-tract crosswalk, with the date closest available to the last seen date of an address to account for the fact that ZIP code definitions may vary and are not nested within Census geographies hud_crosswalk.
 We then distribute the probability to each Census block group within the tracts, again weighting by their relative population shares.
-Success rates for each of these steps are high and detailed in Figure M1.
+Success rates for each of these steps are high and detailed in Figure 6.
 We use the results of this mapping process to create an auxiliary matrix $\mathcal{G}$ of dimensions $N_{\text{ADDRESSES}}\times N_{\text{CBGs}}$.
 Each row of $\mathcal{G}$ defines a probability distribution over Census block groups: an address $i$ belongs to CBG $j$ with probability $\mathcal{G}_{ij}$.
-If the address has been precisely mapped (i.e., geocoded via the Census or ArcGIS geocoders), the corresponding row of $\mathcal{G}$ has a single entry equal to $1$, with other entries 00.
+If the address has been precisely mapped (i.e., geocoded via the Census or ArcGIS geocoders), the corresponding row of $\mathcal{G}$ has a single entry equal to $1$, with other entries 0.
 We use the matrix $\mathcal{G}$ to compute the CBG-to-CBG matrix $E^{(t)}$ via the following equation:
@@
 The second term in this equation maps stayers to the appropriate diagonal entries of the migration matrix: for example, a precisely mapped stayer in CBG $i$ will increment $E^{(t)}_{ii}$ by 1, and someone who stays within ZIP code $a$ will distribute mass (summing to 1) evenly across the diagonal entries corresponding to CBGs within that ZIP code.
-Table M1 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline.
+Summarizing the Processed Infutor Data
+Table 2 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline.
 The number of individuals with active records in the data represents the number of individuals with an interval of activity containing each year at the start of the data processing; the number of CBG moves corresponds to the sum of off-diagonal entries of each matrix $E^{(t+1)}$ and measures the “effective size” of each year’s data: how many movers we actually account for at the beginning of the estimation procedure.
@@
 2018 177,629,604 326,838,199 13,505,890
-Table M1: Yearly Breakdown of Data Summaries.
+Table 2: Yearly Breakdown of Data Summaries.
 Population and addresses accounted by the Infutor and Census datasets during the analysis period (2010-2019).
 (1) Individuals have active records in a given year if the year falls within their interval of activity in the dataset.
-The number of active individuals decreases in recent years, as Infutor records tend to appear when a move is reported.
 Comparison of these values to (2) the Census population shows that Infutor under-counts the population at every given year.
-(3) The number of CBG-to-CBG moves parsed in our dataset.
+(3) The number of moves between Census Block Groups (CBGs) parsed in our dataset.
 Note that an individual with two addresses will have at most one move, some addresses do not constitute residences, some individuals may have a single address in the year thus not move, etc.
+The number of active records drops in recent years because individuals are only marked as “active” when they have a recorded move; many people simply haven’t had a recent move, so they no longer appear as active even though they remain in the population.
+This means that Infutor is more likely to underrepresent non-movers in recent years.
+We employ several measures to mitigate this effect: we pad each individual’s reconciled end date by one year when constructing their active interval, and also rescale Infutor data to match rates of non-movers in Census data.
+We also verify that the decrease in Infutor active records in recent years, which is simply an expected consequence of how active records are defined, does not reflect a broader and more worrisome decrease in Infutor data quality over time.
+To do this, we examine variation over time in the quality metrics discussed in the validations section — the RMSE and correlation between Infutor and Census data for population counts, mover counts, and in-migration rates.
+Results are shown in Table 3.
+While some quality metrics improve slightly over time, and others worsen slightly, overall we do not find consistent evidence that processed Infutor data quality decreases over time.
+Quantity Dataset 2011 2012 2013 2014 2015 2016 2017 2018 2019
+CBG Population Infutor - - - 507 519 531 556 581 -
+MIGRATE - - - 80 91 91 91 88 -
+State-to-State Movers Infutor 3,031 3,684 3,549 3,372 3,607 3,315 2,801 2,518 2,431
+MIGRATE 365 403 400 399 434 410 343 361 383
+State In-Migration Rate Infutor 12.4 14.9 14.4 13.4 14.1 12.4 9.77 6.78 5.56
+(per 1,0001{,}000) MIGRATE 1.68 1.73 1.92 1.48 1.29 1.18 1.07 1.03 1.12
+(a) Root mean squared error (RMSE) for Infutor and MIGRATE, compared to Census data.
+Quantity Dataset 2011 2012 2013 2014 2015 2016 2017 2018 2019
+CBG Population Infutor - - - 0.824 0.826 0.827 0.822 0.818 -
+MIGRATE - - - 0.996 0.995 0.995 0.996 0.996 -
+State-to-State Movers Infutor 0.950 0.956 0.953 0.951 0.954 0.950 0.948 0.936 0.919
+MIGRATE 0.998 0.997 0.997 0.998 0.997 0.998 0.998 0.998 0.998
+State In-Migration Rate Infutor 0.895 0.916 0.911 0.884 0.885 0.896 0.871 0.863 0.809
+MIGRATE 0.980 0.981 0.976 0.987 0.991 0.992 0.993 0.994 0.992
+(b) Pearson correlation for Infutor and MIGRATE, compared to Census data.
+Table 3: Validation of Infutor and MIGRATE estimates over time.
+Year corresponds to year of the Census data release.
+5-Year estimates of Census Block Group Populations are only validated against matrices from 2014 to 2019, which span the full 5-year period.
+Metrics for in-migration rates are weighted by population area.
+More details on computation of these metrics can be found on our validations section.
+Our harmonization process further improves the temporal stability of quality metrics, as can be seen in the MIGRATE rows of Table 3; the correlations between MIGRATE and Census data remain stable and high for all years.
 M.2 Processing Census Data
@@
 Geographic boundaries are not necessarily consistent over time, and ensuring that population counts reflect the same geographic area across multiple years is essential when working with longitudinal data.
-For example, to produce multi-year estimates, the American Community Survey maps every reported address to the corresponding geography in the current year [61].
+For example, to produce multi-year estimates, the American Community Survey maps every reported address to the corresponding geography in the current year acs_methodology.
 We use the 2010 Census boundaries for all addresses to maintain geographic consistency.
@@
 With this approach, we can aggregate population statistics from the fine-grained areas into the coarse areas.
-There was one such county change in 2010-2019 affecting Bedford County in Virginia [62] and a few Census tract or Census block group changes documented yearly in the Census Bureau website (e.g.
-[63] for 2010 to 2011).
-When we aggregate ACS estimates that contain margins of error, we also aggregate margins of error in the L2 norm—as proposed by the ACS [64].
+There was one such county change in 2010-2019 affecting Bedford County in Virginia county_geography_changes and a few Census tract or Census block group changes documented yearly in the Census Bureau website (e.g.
+2011_geography_changes for 2010 to 2011).
+When we aggregate ACS estimates that contain margins of error, we also aggregate margins of error in the L2 norm—as proposed by the ACS acs_userguide.
 For our purposes, area renames are irrelevant.
@@
 Selecting Census datasets
+We rely primarily on three Census datasets: the ACS 5-year CBG populations; ACS 1-year state-to-state flows (from which we use raw flows as well as counts of non-movers and movers); and PEP 1-year county populations.
+PEP makes use of administrative records on births, deaths, and net migration to directly estimate county-level populations; the numbers reported contain the population estimate for July 1st of every year.
+ACS estimates, on the other hand, are a product of year-round surveying aggregated at the correspondent time scale, then released with associated margins of errors that define $90\%$ confidence intervals acs_userguide.
+Our dataset selection and methodology accounts for these different precision levels across the board.
+We would like to avoid matching to datasets that were too noisily estimated.
+To decide which ACS datasets to use in our harmonization process, we examine their level of sampling error.
+We quantify the sampling error by the coefficient of variation (CV): the ratio of the standard deviation to the estimate value.
+Table 4 reports the mean coefficient of variation in ACS estimates.
+In general, the number of non-movers has very low CVs (mean below 2% for counties and below 1% for states); population estimates have considerably lower CVs than flow estimates; state-to-state flows have relatively high CVs (mean around 45% every year), and county-to-county flows have extremely high CVs (mean almost 90%).
+We opt not to match to county-county flows due to the imprecision of raw flow estimates, although we do use them as a validation.
+In-migration rates have much tighter CVs than the corresponding flows (mean around 15% and 5% for counties and states, respectively), and we also use those as validations.
 2010 2011 2012 2013 2014 2015 2016 2017 2018 2019
@@
 State In-Migration Rate - 4.95 4.68 4.79 4.65 4.70 4.76 4.72 4.85 5.02
-Table M2: ACS Coefficients of Variation.
-Average coefficient of variation (in %) of the non-zero estimates in each dataset.
-CVs were computed as the ratio of the standard error (obtained by dividing the margin of error by $1.645$) by the estimate [64].
+Table 4: Uncertainty in Census datasets.
+Average coefficient of variation (CV) of the non-zero estimates in each American Community Survey (ACS) dataset (in %).
+CVs were computed as the ratio of the standard error (obtained by dividing the margin of error by $1.645$) by the estimate acs_userguide.
 We did not use flows released in 2010 (as they would report moves from 2009).
-We rely primarily on three Census datasets: the ACS 5-year CBG populations; ACS 1-year state-to-state flows (from which we use raw flows as well as counts of non-movers and movers); and PEP 1-year county populations.
-PEP makes use of administrative records on births, deaths, and net migration to directly estimate county-level populations; the numbers reported contain the population estimate for July 1st of every year.
-ACS estimates, on the other hand, are a product of year-round surveying aggregated at the correspondent time scale, then released with associated margins of errors that define $90\%$ confidence intervals [64].
-Our dataset selection and methodology accounts for these different precision levels across the board.
-We would like to avoid matching to datasets that were too noisily estimated.
-To decide which ACS datasets to use in our harmonization process, we examine their level of sampling error.
-We quantify the sampling error by the coefficient of variation (CV): the ratio of the standard deviation to the estimate value.
-Table M2 reports the mean coefficient of variation in ACS estimates.
-In general, the number of non-movers has very low CVs (mean below 2% for counties and below 1% for states); population estimates have considerably lower CVs than flow estimates; state-to-state flows have relatively high CVs (mean around 45% every year), and county-to-county flows have extremely high CVs (mean almost 90%).
-We opt not to match to county-county flows due to the imprecision of raw flow estimates, although we do use them as a validation.
-In-migration rates have much tighter CVs than the corresponding flows (mean around 15% and 5% for counties and states, respectively), and we also use those as validations.
 We perform a synthetic simulation study to verify that the datasets we do match to — state-to-state flows, and CBG population estimates — have sufficiently low sampling error to be reliably used.
@@
 Based on these analyses, we harmonize the entries of each yearly matrix $E^{(t)}$ with the following data: (1) CBG populations (both from 2010 Census and 5-year ACS); (2) ACS 1-year counts of movers and non-movers by state; (3) ACS 1-year state-to-state flows; and (4) PEP 1-year county populations.
+Population count data at the sub-county level was obtained via IPUMS National Historical Geographic Information System, along with socieoconomic and demographic data used for our bias analyses ipums_nhgis; population flows and movers data was obtained directly via ACS acs_migration_flows; population count data at county level was obtained via the Population Estimates Program PEP.
 Accounting for natural population increase and international migration
 Using Census datasets to constrain MIGRATE estimates requires ensuring that the Census datasets reflect the same population accounted for in the Infutor matrix $E^{(t)}$.
-Each entry $E^{(t)}_{ij}$ represents the expected number of people alive with a residence in area $i$ at year $t-1$ and with a residence in area $j$ at year $t$, and diagonal entries $E^{(t)}_{ii}$ also include individuals who resided in area $i$ at year $t-1$ but died or emigrated.
+Each entry $E^{(t)}_{ij}$ represents the expected number of people alive with a residence in area $i$ at year t1t-1 and with a residence in area $j$ at year $t$, and diagonal entries $E^{(t)}_{ii}$ also include individuals who resided in area $i$ at year t1t-1 but died or emigrated.
 We process the Census datasets so they reflect the same population, by accounting for changes in the populations at year $t$ due to births, deaths, and international migration.
-Specifically, when using Census populations from year $t$ to scale entries of $E^{(t)}$, we want to remove from each Census area population the natural increase in population (i.e., births minus deaths) and the net international migration (i.e., immigrants minus emigrants); when using Census flows from year $t-1$ to year $t$, we want to add counts of deaths and emigrants back into the diagonal as non-movers.
+Specifically, when using Census populations from year $t$ to scale entries of $E^{(t)}$, we want to remove from each Census area population the natural increase in population (i.e., births minus deaths) and the net international migration (i.e., immigrants minus emigrants); when using Census flows from year t1t-1 to year $t$, we want to add counts of deaths and emigrants back into the diagonal as non-movers.
 We adjust the Census constraints individually for every matrix $E^{(t)}$.
 To do this adjustment, we make use of the PEP components of population change and ACS international immigration estimates.
-PEP releases estimates for the number of births, deaths, net international migration, and net domestic migration yearly at both a county and a state level [65].
-To build yearly population estimates, PEP directly adds to the previous year’s estimates the natural increase, net international migration, and net domestic migration [^8] PEP produces estimates that are geographically consistent: county-level estimates must aggregate precisely to state-level estimates, which must aggregate to nation-level estimates [65].
+PEP releases estimates for the number of births, deaths, net international migration, and net domestic migration yearly at both a county and a state level PEP.
+To build yearly population estimates, PEP directly adds to the previous year’s estimates the natural increase, net international migration, and net domestic migration [^7] PEP produces estimates that are geographically consistent: county-level estimates must aggregate precisely to state-level estimates, which must aggregate to nation-level estimates PEP.
 To achieve such precision, they also estimate small residuals at the county and state levels.
@@
 We do not scale any entries to zero because the Census zeroes are noisily estimated, and subsequent multiplicative re-scalings cannot correct erroneous zero scalings.
-We detail below, in order, the scalings we conduct.
+We detail below, in order, the scalings we perform.
 Harmonizing with CBG population data
@@
 For example, when we take the average sum of each row across the five matrices $\{E^{(2011)},E^{(2012)},E^{(2013)},E^{(2014)},E^{(2015)}\}$, this should be equivalent to the population for the corresponding CBG in the 2011-2015 ACS.
-Imposing these constraints corresponds to solving a linear system for each row using non-linear least squares (to impose the constraint that populations cannot be negative).
-Harmonizing with yearly state non-movers data
+We also impose a boundary constraint to match the 2010 population to that reported by the Decennial Census.
+Imposing these constraints corresponds to solving the non-negative least squares optimization problem
+minimize $\|A\cdot\vec{x}-\vec{b}\|$
+s.t. $\vec{x}\geq 0$ (2)
+where
+[…]
+for each row (i.e.
+each CBG).
+The vector $\vec{x}$ represents the estimated population for the CBG in each year from 2009 to 2019, and we impose the constraint that $\vec{x}\geq 0$ to ensure populations are non-negative.
+We use these estimated CBG populations to harmonize our flow matrices, by scaling each row to match the estimated CBG population in the relevant year.
+Harmonizing with yearly data on state movers and non-movers
 Next, we harmonize $E^{(t)}$ with the counts of movers and non-movers by state from 1-year ACS data.
@@
 We also aggregate the diagonal entries of columns corresponding to each state and scale them to match the count of non-movers by state.
-These scalings (which treat the population remaining within each CBG as equivalent to the population of non-movers) assume that very few people move to another location within the same CBG, an assumption substantiated by previous research.
-[^9] Research has shown that median distance people moved in the US was about 10 to 15 miles in the 2010-19 period, which is far more than the average 2010 CBG radius (1.7 miles) and close to the 98-th percentile of radii [66].
+These scalings (which treat the population remaining within each CBG as equivalent to the population of non-movers) assume that very few people move to another location within the same CBG, an assumption substantiated by previous research: the median distance people moved in the US was about 10 to 15 miles in the 2010-19 period, which is far more than the average 2010 CBG radius (1.7 miles) and close to the 98th percentile of CBG radii lautz_long-distance_2023.
+Mathematically, let $S^{(t)}_{k}$ be population of state $k$ in year $t$ and $R^{(t)}_{k}$ the population of state $k$ in year $t$ which did not move in the previous year.
+Then we multiply every diagonal entry $E^{(t)}_{xx}$ where CBG $x$ lies in state $r$ by
+$\dfrac{R^{(t)}_{r}}{{\sum_{i}E_{ii}^{(t)}\cdot\mathds{1}\{\text{CBG }i\;\in\;\text{state }r\}}}$ (3)
+and then multiply every off-diagonal entry $E^{(t)}_{xy}$ ($x\neq y$) where CBG $x$ lies in state $r$ and CBG $y$ lies in state $s$ by
+$\dfrac{S^{(t)}_{s}-R^{(t)}_{s}}{{\sum_{i,j\;:\;i\neq j}E_{ij}^{(t)}\cdot\mathds{1}\{\text{CBG }j\;\in\;\text{state }s\}}}$ (4)
+[…]
 Harmonizing with yearly state-to-state flow data
@@
 We do not scale to match zero flows.
+Mathematically, let $F^{(t)}_{rs}$ be the number of movers between states $r$ and $s$ from year t1t-1 to year $t$.
+Then we multiply every entry $E^{(t)}_{xy}$ where CBG $x$ lies in state $r$ and CBG $y$ lies in state $s$ by
+$\dfrac{F^{(t)}_{rs}}{\sum_{i,j}E^{(t)}_{ij}\cdot\mathds{1}\{\text{CBG }i\;\in\;\text{ state }r\}\cdot\mathds{1}\{\text{CBG }j\;\in\;\text{ state }s\}}$ (5)
+whenever $F^{(t)}_{rs}\neq 0$. We scale both people who move between states ($r\neq s$) and who remain within the same state ($r=s$).
 Harmonizing with yearly county population data
-We finally harmonize $E^{(t)}$ with county populations from 1-year PEP data.
+Finally, we harmonize $E^{(t)}$ with county populations from 1-year PEP data.
 We choose to match $E^{(t)}$ to PEP populations last due to their superior reliability.
-In addition, PEP estimates are used internally by the Census Bureau as controls for many other datasets [67], and ensuring MIGRATE estimates agree with PEP can lead to better downstream agreement in other datasets.
-To match to PEP county populations, we apply an IPF-based algorithm, alternating between two steps until convergence: (1) we aggregate the entries of $E^{(t)}$ by columns corresponding to each county and scale them to match the PEP county population at year $t$ and (2) we aggregate the entries of $E^{(t)}$ by rows corresponding to each county and scale them to match the PEP county population at year $t-1$.
-Our algorithm runs for $3,000$ iterations, which we verify is sufficient for convergence.
-Specifically, at the end of 3,000 iterations, the total update to the matrix $E^{(t)}$—that is, the $L^{1}$ distance between the resulting matrices in two subsequent iterations—falls to the order of $10^{-7}$.
-[…]
-Supplementary information Supplementary Information is available detailing additional methods, validations, and findings.
-Declarations
-[…]
-Data availability MIGRATE is available upon request for non-profit research use at migrate.tech.cornell.edu.
-[…]
-Code availability Code to reproduce our research findings is available on GitHub: https://github.com/gsagostini/MIGRATE.
+In addition, PEP estimates are used internally by the Census Bureau as controls for many other datasets us_census_bureau_population_2022, and ensuring MIGRATE estimates agree with PEP can lead to better downstream agreement in other datasets.
+To match to PEP county populations, we apply a classical IPF-based algorithm, alternating between two steps until convergence: (1) we aggregate the entries of $E^{(t)}$ by columns corresponding to each county and scale them to match the PEP county population at year $t$ and (2) we aggregate the entries of $E^{(t)}$ by rows corresponding to each county and scale them to match the PEP county population at year t1t-1.
+More formally, let $P^{(t)}_{c}$ be the population of a given county $c$ at a given year $t$.
+We start with the matrix $M^{0}:=E^{(t)}$.
+We then perform the following updates for iterations $n=1,2,\dots N$.
+For odd $n$, we scale the blocks of columns of the matrix to match the county populations in year $t$.
+Specifically, for each CBG $y$ lying within county $q$, we multiply entries $M^{n-1}_{xy}$ by the scaling factor
+$\dfrac{P^{(t)}_{q}}{\sum_{i,j}M^{n-1}_{ij}\cdot\mathds{1}\{\text{CBG }j\in\text{county }q\}}$ (6)
+For even $n$, we scale blocks of rows of the matrix to match the county populations in year t1t-1.
+Specifically, for each CBG $x$ lying within county $p$, we multiply entries $M^{n-1}_{xy}$ by the scaling factor
+$\dfrac{P^{(t-1)}_{p}}{\sum_{i,j}M^{n-1}_{ij}\cdot\mathds{1}\{\text{CBG }i\in\text{county }p\}}$ (7)
+Our algorithm runs for $6,000$ iterations, which we verify is sufficient for convergence. Specifically, we track the L1 distance between the resulting matrices in two subsequent iterations.
 Appendix A Additional details on validations
 Here we provide additional details on MIGRATE validations.
-§A.1 provides additional validation metrics that complement Figure 2; §A.2 presents additional details on our held-out data validations; §A.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §A.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
-A.1 Additional validation metrics
-Fig.
-S1: Detailed flow validations of MIGRATE.
-To complement the findings shown in Figure 2, we present validations with respect to area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e.
-the off-diagonal entries of the migration matrix), and raw in-migration (i.e.
-the number of people arriving at an area).
-We show (a-c) Pearson correlations of estimates and selected ACS flows as well as (d-f) how much MIGRATE reduces RMSE compared to Infutor.
-Figure S1 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e.
-the off-diagonal entries of the migration matrix), and raw in-migration (i.e.
-the number of people arriving at an area).
-We report these metrics at both the state and county level; state-to-state flows were used in the estimation procedure, county-to-county flows were held out.
-For all datasets, MIGRATE estimates are highly correlated with Census estimates ($\rho$ above $0.957$ in all cases) and reduce RMSE relative to the Infutor data.
-A.2 Validating the estimation procedure using held-out data
+We include additional details on our held-out data validations; a semi-synthetic data validation where we verify our method reconstructs a known ground-truth matrix; details on related work documenting biases of Infutor, which is the subject of our Figure 3; and additional socioeconomic and demographic dimensions on which we assessed biases, including undocumented immigrants.
+A.1 Validating the estimation procedure using held-out data
 MIGRATE CBG Held-Out State Held-Out
@@
 Pearson correlations with MIGRATE, which is estimated using all data sources, are presented for comparison.
+As noted in the main text, perfect correlations between MIGRATE and all flows or stayers are partially driven by variation in area size and the predominance of non-mover counts among flows.
 MIGRATE CBG Held-Out State Held-Out
@@
 Values are average reductions across all years with overlapping ground-truth data releases.
-Table S1: Validation of estimates in held-out data settings.
+Supplementary Table 1: Validation of estimates in held-out data settings.
 The full model validations are repeated for comparison.
@@
 $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates).
-Table S1 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor.
-Values for the original MIGRATE estimates, reported in Figure 2 and Figure S1, are reproduced for easy comparison.
+Supplementary Table 1 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor.
+Values for the original MIGRATE estimates, reported in Figure 2, are reproduced for easy comparison.
 The CBG held-out and State held-out estimates remain highly correlated with all Census datasets ($\rho$ above $0.629$); they also reduce RMSE relative to Infutor.
@@
 Together, these results illustrate both that (a) our estimation pipeline generalizes well to held-out data and that (b) all datasets used in the training help improve concordance with Census data, justifying their inclusion.
+A.2 Validating the harmonization procedure with semi-synthetic data
+We further validate our harmonization procedure using semi-synthetic data, a standard check for statistical methods when ground-truth data is lacking balachandar2024domain, agostini2024bayesian, pierson2019inferring.
+Specifically, we assess how well our harmonization procedure can recover a known CBG-to-CBG flow matrix, given only a perturbed version of that matrix and the same set of marginal constraints available in our main analysis (e.g., CBG-level populations).
+To generate realistic semi-synthetic data, we use the MIGRATE matrices as the ground-truth CBG-to-CBG flow matrices $M$, since these are likely more realistic than purely synthetic data generated from stylized models.
+These ground-truth matrices $M$ are not used as input to our harmonization method, but only as validation.
+From $M$, we produce the perturbed CBG-to-CBG flow matrix $E$ by adding multiplicative noise; we describe this process further below.
+We also compute the same marginals of $M$ that are available in our main analysis: specifically, CBG populations; state-level counts of movers and non-movers; state-to-state flows; and county populations.
+We use these marginals and the perturbed matrix $E$ as inputs to our harmonization method.
+We assess how well our method recovers the ground-truth matrix $M$ using the same metrics described in the main text (reduction in RMSE relative to using $E$ alone, and Pearson correlation).
+To thoroughly test the robustness of our method, we consider two processes for generating $E$ from $M$. We now describe these processes.
+Harmonizing in the presence of structured multiplicative noise
+Our harmonization procedure assumes that the difference between the ground-truth matrix $M$ and the perturbed matrix $E$ can be described by a series of multiplicative scalings with a given structure (e.g., CBG-level scalings on each row) as described in the Methods Section.
+We thus first test how well our harmonization procedure recovers the ground-truth matrix if the perturbation does indeed match that structure.
+Specifically, we draw independent Log-Normal random variables that multiply (1) each row of $M$ (mirrors CBG population fitting), (2) each group of diagonal entries within the same state (mirrors state non-movers fitting), (3) each group of off-diagonal entries in columns within the same state (mirrors state movers fitting), (4) each group of entries within the same state-to-state pair (mirrors state flows fitting) and (5) each group of rows within the same county, and each group of columns within the same county (mirrors county population fitting).
+We draw the Log-Normal noise using standard deviations $\tau$ ranging from $0.05$ to $0.20$.
+We choose the upper bound on this range because it exceeds the level of variation we observe in real-world data.
+Specifically, across states (Figure 3a), the biases between ground-truth and Infutor populations have standard deviation 0.11; across counties, the standard deviation is 0.19, both exceeded by our upper bound of $\tau=0.20$.
+(We do verify, however, that results remain robust up to much larger $\tau$ values of 1.00).
+We find that our harmonization method performs very strongly in this setting.
+For all standard deviations of noise, our inferred matrix achieves Pearson correlations above 0.99 either on the full CBG-CBG matrix or when restricting to CBG-CBG movers.
+The harmonized matrix also reduces RMSE (relative to the original perturbed matrix $E$) by an average of 97.7% across all years and noise levels (sd 0.06%), and by 91.9% when restricting to off-diagonal entries (sd 0.67%).
+These RMSE reductions remain consistently strong across the range of noise scales.
+The small residual error is an expected result of the optimization procedure; for example, we only fit once to CBG-level marginals to avoid overfitting, so are not guaranteed to match these marginals.
+Harmonizing in the presence of real-world bias and independent noise
+We next assess the performance of our harmonization procedure when the perturbation does not match the structure it assumes, but consists both of real-world bias and independent noise.
+Specifically, we produce the perturbed matrix $E$ from $M$ as follows:
+$E_{ij}=M_{ij}\cdot\exp\left\{b\cdot\left(w_{i}+w_{j}\right)+\sigma\cdot Z_{ij}\right\}$
+where $w_{i}$ is the z-scored share of the white population in CBG $i$ (capturing real-world biases, as shown in Figure 3c) and $Z_{ij}$ are i.i.d.
+standard normal variables.
+The parameters $b$ and $\sigma$ control the relative scales of the bias and the independent noise, respectively.
+We assess performance across a range of $b$ and $\sigma$, choosing $b$ to represent a realistic level of real-world bias based on our results in Figure 3c, and noise scales $\sigma$ up to 0.20, as above.
+We would expect high levels of independent noise to make it very challenging to recover the off-diagonal entries of the ground-truth matrix.
+This is because the diagonal entries of the migration matrix are much bigger than the off-diagonal entries (since most people do not move); thus, the scaling factors learned by the harmonization procedure will be dominated by the diagonal entries.
+This is generally reasonable: if we want to estimate bias in the movers from a CBG, the bias of the entire CBG population likely serves as a good proxy.
+However, this logic fails under independent noise, since noise in the diagonals provides no information about noise in the off-diagonals.
+Importantly, the challenge of correcting for independent noise applies not just to our harmonization procedure, but to any correction procedure, since correcting such noise would require information about the ground truth for each entry.
+Supplementary Fig.
+1: Experimental validations of our method.
+Pearson correlation and reduction in RMSE achieved by our harmonization procedure on semi-synthetic data with real-world bias and independent noise.
+We report average metrics across all years; error bars plot standard deviation across years (from 2010-11 to 2018-19, $n=9$).
+Circles represent metrics computed on all entries of the flow matrix; squares represent metrics computed on movers only (off-diagonal elements).
+Columns assess metrics for flows at the CBG level, tract level, county level, and state level.
+We assess results when varying the level of bias (x-axis) and independent noise (line color).
+Supplementary Figure 1 plots the Pearson correlation and reduction in RMSE for flows at every geographic scale.
+We compute metrics with respect to all flows and movers only (off-diagonal entries) for varying levels of noise and bias.
+Results are consistent with the analysis above.
+When considering all flows (as opposed to just off-diagonal entries), error is reduced almost completely and correlations are nearly perfect at all geographic scales (circular points).
+When restricting to off-diagonal entries specifically (square points) on the more granular geographic scales (tract and CBG levels), we find the largest error reductions when the scale of the independent noise is not too large relative to the bias, as expected.
+Pearson correlations decrease as bias and noise increase, but remain relatively high (above $0.91$ in all cases).
+State and county-level metrics for movers remain robust at all noise and bias scales because we have better marginal data with which to correct them (including state-to-state movers and county populations).
+Collectively, these synthetic experiments show that our harmonization procedure performs well in recovering the ground-truth matrix as long as the structure of the perturbation lies within the family it is able to estimate; in contrast, purely independent noise is unsurprisingly harder to correct (both for our procedure and for any correction procedure).
+The family of perturbations our harmonization procedure can estimate is highly flexible (due to the number of scalings of varying shapes it incorporates) and indeed more flexible than classic iterative proportional fitting, which is widely used in similar demographic applications birkin_synthesissynthetic_1988, wong_reliability_1992, simpson_combining_2005, chang2021mobility, chang2024inferring and estimates only row and column-level scalings.
+The assumption that the perturbation is structured, as opposed to entirely independent for each entry, is also substantiated by our and previous findings of systematic demographic bias in Infutor data, which would not emerge under purely independent noise.
 A.3 Related work documenting biases in the Infutor data
@@
 We include here a non-exhaustive review of works with documentation of Infutor biases.
-As we report in Table M1, Infutor undercounts the national US population.
-This is consistent with prior work; a similar finding is reported by Qian and Tan [21], Bernstein et al.
-[19], and Diamond et al.
-[17] among others.
-Phillips [18] additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers.
-A contrasting finding, reported by Diamond et al.
-[4], is that Infutor might overcount the population according to the 2000 Census in San Francisco.
+As we report in Table 2, Infutor undercounts the national US population.
+This is consistent with prior work; a similar finding is reported by qian2021effects, bernstein2022contribution, and diamond_effects_2019 among others.
+phillips_measuring_2020 additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers.
+A contrasting finding, reported by diamond_effect_2020, is that Infutor might overcount the population according to the 2000 Census in San Francisco.
 Such a finding might be driven by specific spatiotemporal idiosyncrasies of the San Francisco data in 2000, which is not included in the subset we analyze; however, the bulk of the literature, and all the literature which analyzes the same subset of the data we do, is consistent with our finding that Infutor undercounts the data.
-Diamond et al.
-[4] also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
-In Figure 3, we also assess whether under- or over-represents demographic subpopulations.
-(Prior to conducting this assessment, we scale the Infutor population to match the overall population, so any discrepancies are not caused merely by overall undercounting.) Ramiller et al.
-[22] similarly study representation biases of Infutor in King County, Washington, during the 2015-2019 5-year period.
+diamond_effect_2020 also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
+In Figure 3, we also assess whether under- or over-represents demographic subpopulations [^8] Prior to conducting this assessment, we scale the Infutor population to match the overall population, so any discrepancies are not caused merely by overall undercounting.
+ramiller_residential_2024 similarly study representation biases of Infutor in King County, Washington, during the 2015-2019 5-year period.
 They find that Infutor undercounts the county population while overrepresenting White, older, home-owning, and high-income populations.
@@
 A.4 Assessment of additional demographic biases
-Fig. S2: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
-We documented biases in the Infutor dataset geographically (Figure 3a) and on the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c).
-Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure S2).
+Supplementary Fig.
+2: Additional analysis of socioeconomic and demographic biases at CBG population level.
+Variables are taken from the 5-year ACS CBG estimates.
+Bias is averaged over $n=5$ population data releases (American Community Survey 5-year estimates, 2015 through 2018) Error bars represent standard deviations across these releases.
+We documented biases in the Infutor dataset geographically (Figure 3a) and along the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c).
+Here, we provide additional analyses of demographic bias by sex, educational attainment, number of children per household, and country of birth (Supplementary Figure 2).
 We find that Infutor displays negligible bias on the basis of sex.
@@
 This finding is consistent with the Infutor data documentation, which describes a lack of records for people under the age of 18 (therefore, children will be undercounted).
+Infutor also undercounts foreign-born residents by about 12%.
 MIGRATE essentially eliminates biases along all these demographic dimensions.
+Biases in Representation of Immigrant Populations
+Because of our finding that Infutor undercounts the foreign-born population, and previous research hypothesizing that Infutor may undercount undocumented immigrants specifically downes2023can, bernstein2022contribution, phillips_measuring_2020, we conduct further analyses investigating Infutor’s representation of undocumented immigrants.
+Infutor draws from many data sources — including cell phone bills, phone books, credit header files, public government records, property deeds, county property records, vehicle warranties, and data from vehicle repair and maintenance providers diamond_effects_2019, phillips2024homelessness, boar2023consumption — and some of these sources may contain data for undocumented immigrants.
+There were an estimated 11 million undocumented immigrants in the United States in the 2018-2022 period, about 27% of the estimated foreign-born population baker_estimates_2024.
+To measure the biases of Infutor in representing the undocumented immigrant population, we investigate whether the relative population error correlates with estimates of the undocumented population.
+We use data on the yearly undocumented population per state provided by the Pew Research Center passel_what_2024, which is the most granular data available.
+As expected from the results in Figure 3a, which documents bias in Infutor data at the county level, there is geographic bias in the Infutor data at the state level as well.
+Infutor undercounts the population in states with larger proportions of undocumented immigrants: there is a Spearman correlation of $\rho=-0.401$ between the state-level Infutor error and the proportion of undocumented immigrants per state, averaging across the decade (Supplementary Figure 3b).
+Supplementary Fig.
+3: Infutor undercounts the undocumented immigrant population.
+We study the representation of the undocumented immigrant population by Infutor at the state level.
+(a) Average errors in state populations in Infutor data relative to Census data.
+MIGRATE estimates remove all state-level errors by construction.
+(b) Spearman correlation between the estimated share of undocumented immigrants in the state population (x-axis) and error in Infutor population estimates (y-axis).
+Infutor is more likely to undercount states with larger proportions of undocumented immigrants.
+Importantly, as with other demographic biases in the Infutor dataset, harmonizing the dataset with Census data reduces biases with respect to immigrant populations as well.
+Census statistics do include undocumented immigrants us_census_bureau_about_2025, and so rescaling to match Census populations will mitigate Infutor’s biases.
+[^9] However, Census data may also be imperfect with respect to its coverage of undocumented migrants if they are, for example, less likely to respond to surveys.
+In particular, MIGRATE corrects all bias at state level by construction. As shown in Supplementary Figure 2, MIGRATE also significantly reduces the bias for the overall foreign-born population.
 Appendix B Additional analyses of national migration
-In this section we provide additional analyses of national migration trends discussed in Section 2.3.
+In this section we provide additional analyses of national migration trends to complement the findings of Section 4.
 B.1 Out-migration rates by CBG type
@@
 Top 128 11,851,831
-Table S2: Out-migration by group.
+Supplementary Table 1: Stratified out-migration.
 Out-migration rates (per 1,000 people) and raw counts of out-movers categorized by plurality race, urbanization level, and income quartile.
 We average the values across all years in the period.
-Table S2 reports the average out-migration rates for each of the ten CBG groups we analyze in the main text.
+Supplementary Table 1 reports the average out-migration rates for each of the ten CBG groups we analyze in the main text.
 The average out-migration rate in all CBGs is $13.0\%$, with $0.9\%$ standard deviation across years.
@@
 B.2 Additional statistics on migration homophily
-Fig.
-S3: Additional national migration statistics.
-(a) Flows between the ten types of CBGs discussed in the main text relative to the share of all movers who moved to CBGs in the particular group.
-Each cell reports the ratio share of movers from origin CBG group moving to destination CBG groupshare of all movers moving to destination CBG $\frac{\text{share of movers from origin CBG group moving to destination CBG group}}{\text{share of all movers moving to destination CBG group}}$; for example, the top left value of $\times1$ reports the ratio of the share of movers from plurality white CBGs who move to plurality white CBGs (90%) to the share of movers from all CBGs who move to plurality white CBGs (80%).
-(b) Flows between the ten types of CBGs discussed in the main text when restricting to out-of-county movers demonstrate that homophily persists for long-distance moves.
-In Figure S3, we extend the analysis of national homophily patterns in migration documented in Figure 4a.
-Figure S3a shows the shares of movers moving between groups of CBGs relative to the share of all movers who move to that group of CBGs (that is, a version of Figure 4a where the value in each cell is divided by the corresponding value in the row ‘Movers from all CBGs’).
-Homophily in migration might occur because many moves are local (within county or within 5 miles) and demographics are spatially correlated.
-To investigate this possibility, we restrict our analysis to out-of-county moves in Figure S3b.
-Absolute racial homophily is less prominent—i.e.
-there are smaller shares in the diagonal of the race by race submatrix—yet still present: out-of-county movers from plurality Black, Asian, and Hispanic CBGs are 4.3$\times$, 5.5$\times$, and 3.1$\times$ likelier than out-of-county movers as a whole to move to CBGs with the same plurality race group (vs.
-5.5$\times$, 14.6$\times$, and 4.3$\times$ for all movers).
+[…]
+In Supplementary Figure 1, we extend the analysis of national homophily patterns in migration documented in Figure 4a.
+Supplementary Figure 1a shows the shares of movers moving between groups of CBGs relative to the share of all movers who move to that group of CBGs (that is, a version of Figure 4a where the value in each cell is divided by the corresponding value in the row ‘Movers from all CBGs’).
+[…]
 B.3 Additional statistics on economic mobility in migration
-Fig.
-S4: Further national migration statistics on upwards mobility.
+Supplementary Fig.
+2: Further national migration statistics on upwards mobility.
 Probability of moving to a (a) top-income quartile CBG, (b) bottom-income quartile CBG, and (c) higher-income CBG conditioned on current residence median income and plurality race.
@@
 (d) Movers remain more likely to move to top-income-quartile CBGs than bottom-income-quartile CBGs throughout the decade, with the disparity increasing in magnitude over time.
-Figure S4 provides additional analyses supporting our finding of racial disparities in economic mobility (Figure 4b): movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs, even when controlling for income of origin CBG.
-We show that analogous racial disparities still emerge when looking at probability of moving to a CBG in the top income quartile (Figure S4a) or bottom income quartile (Figure S4b) as opposed to the probability of moving to a higher-income CBG (as in the main text).
-These disparities also emerge when controlling for origin CBG income percentile (Figure S4c), as opposed to decile, as in the main text.
-We also observe that the gap in upward mobility widens during the decade: the share of movers going to a top income quartile CBG increases from 31% to 33% while the share of movers going to a bottom income quartile CBG decreases from 19% to 17%; overall, the ratio of out-movers moving to top income quartile versus bottom income quartile CBGs increases from 1.64 to 1.95.
+Supplementary Figure 2 provides additional analyses supporting our finding of racial disparities in economic mobility (Figure 4b): movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs, even when controlling for income of origin CBG.
+We show that analogous racial disparities still emerge when looking at probability of moving to a CBG in the top income quartile (Supplementary Figure 2a) or bottom income quartile (Supplementary Figure 2b) as opposed to the probability of moving to a higher-income CBG (as in the main text).
+These disparities also emerge when controlling for origin CBG income percentile (Supplementary Figure 2c), as opposed to decile, as in the main text.
+We also observe that the gap in upward mobility widens during the decade: the share of movers going to a top income quartile CBG increases from 31% to 33% while the share of movers going to a bottom income quartile CBG decreases from 19% to 17% (Supplementary Figure 2d); overall, the ratio of out-movers moving to top income quartile versus bottom income quartile CBGs increases from 1.64 to 1.95.
 B.4 Additional statistics on moving distance
-Fig. S5: Further national migration statistics on distance moved. (a) Moving distance by year. (b) Fraction of movers moving within tract, county, and state.
-Figure S5 provides additional statistics on moving distance, expanding the findings in Figure 4c.
-Figure S5a reports how moving distance changes over time, showing a general increase over the decade: the share of movers moving over 50 miles, for example, grows from about 22% to about 26%, and the average moving distance increases from 154 to 191 miles in the decade.
-Figure S5b reports moving distance stratified by geographic boundary (that is, the share of movers remaining within tract, county, and state).
-Findings are generally consistent with those reported in Figure 4c, and also with published ACS figures, which state that over half of moves happen within counties [15].
-Appendix C Case Study: destinations of New York City outmovers
-Fig.
-S6: Average income of destination CBGs and counties for New York City out-movers in the 2010-19 period.
+Supplementary Fig. 3: Further national migration statistics on distance moved. (a) Moving distance by year. (b) Fraction of movers moving within tract, county, and state.
+Supplementary Figure 3 provides additional statistics on moving distance, expanding the findings in Figure 4c.
+Supplementary Figure 3a reports how moving distance changes over time, showing a general increase over the decade: the share of movers moving over 50 miles, for example, grows from about 22% to about 26%, and the average moving distance increases from 154 to 191 miles in the decade.
+Supplementary Figure 3b reports moving distance stratified by geographic boundary (that is, the share of movers remaining within tract, county, and state).
+Findings are generally consistent with those reported in Figure 4c, and also with published ACS figures, which state that over half of moves happen within counties kerns-damore_migration_2023.
+Appendix C Additional case studies of local migration patterns
+In this section we provide two case studies of how MIGRATE reveals policy-relevant local migration patterns absent from publicly available data in New York City: a study of socioeconomic disparities on out-mover destinations, and a description of migration patterns to and from public housing buildings.
+C.1 Destinations of New York City outmovers
+Supplementary Fig.
+1: Average income (in US dollars \$) of destination CBGs and counties for New York City out-movers in the 2010-19 period.
 The large map uses MIGRATE estimates and 5-year ACS CBG data.
@@
 We provide a case study contrasting MIGRATE with publicly available county-to-county migration data when analyzing migration stratified by a demographic or socioeconomic feature.
-Figure S6 plots the average income of the CBG to which a person moves according to their previous CBG of residence.
+Supplementary Figure 1 plots the average income of the CBG to which a person moves according to their previous CBG of residence.
 To compute this quantity for a given CBG $i$, we average the mean income of all destination CBGs $j$ for out-movers of CBG $i$ weighted by the number of out-movers from $i$ to $j$.
@@
 Comparing these results to the data available at county level reveals striking within-county variation.
-For example, the average income of the county to which people from Manhattan moved was \$$\$82,934$—lower than that of the other four counties in New York City.
-MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest average income among all New York City CBGs — above \$$\$175,000$ at times — while movers from other Manhattan CBGs moved to areas with some of the lowest average incomes.
+For example, the average income of the county to which people from Manhattan moved was $\mathdollar 82,934$—lower than that of the other four counties in New York City.
+MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest average income among all New York City CBGs — above $\mathdollar 175,000$ at times — while movers from other Manhattan CBGs moved to areas with some of the lowest average incomes.
 Intra-county socioeconomic differences factor heavily in out-migration: Manhattan out-movers with the highest-income destinations tended to reside in the central and lower parts of the county (which tend to have higher incomes), whereas those with lower-income destinations came mostly from northern CBGs (which tend to have lower incomes).
@@
 Overall, this analysis illustrates the important demographic heterogeneity in migration which is concealed at the county level.
-Appendix D Case Study: New York City Housing Authority
-Ensuring that residents have access to affordable housing is an important and challenging policy goal [68].
-One way of achieving this is through the creation and management of public housing units: for example, the New York City Housing Authority (NYCHA) is the largest public housing authority in the US, providing affordable housing to over 500,000 authorized residents [69].
-We provide a case study illustrating how the MIGRATE dataset can be used to understand which residents make use of public housing units, thereby illuminating the effect of this component of affordable housing policy.
-Fig.
-S7: We use MIGRATE to understand migration to and from New York City Housing Authority (NYCHA) developments.
+C.2 New York City Housing Authority
+Supplementary Fig.
+2: We use MIGRATE to understand migration to and from New York City Housing Authority (NYCHA) developments.
 (a) A map of the 114 Census Block Groups (marked in red) where all residential units are owned by NYCHA (i.e., where all the population lives in a NYCHA development).
@@
 (d) Demographics of next residence CBG for out-movers.
-We identify all 114 New York City Census Block Groups whose residential units are entirely owned by NYCHA using New York City’s Primary Land Use Tax Lot Output (PLUTO) dataset [70] (Fig.
-S7a).
+Ensuring that residents have access to affordable housing is an important and challenging policy goal favilukis2023affordable.
+One way of achieving this is through the creation and management of public housing units: for example, the New York City Housing Authority (NYCHA) is the largest public housing authority in the US, providing affordable housing to over 500,000 authorized residents nycha_factsheet_2024.
+We provide a case study illustrating how the MIGRATE dataset can be used to understand which residents make use of public housing units, thereby illuminating the effect of this component of affordable housing policy.
+We identify all 114 New York City Census Block Groups whose residential units are entirely owned by NYCHA using New York City’s Primary Land Use Tax Lot Output (PLUTO) dataset nyc_pluto (Supplementary Figure 2a).
 We compare migration patterns for these CBGs, which we refer to as “NYCHA CBGs”, to migration patterns of two other groups of non-NYCHA CBGs: (i) nearby CBGs (within 250 meters of a NYCHA development) and (ii) all non-NYCHA CBGs in New York City.
 [^10] We excluded from the analysis all 345 CBGs (about 5.3% of all New York City CBGs) where some but not all of the residential units are owned by NCYHA.
-NYCHA CBGs have had consistently lower in-migration and out-migration rates than the two comparison groups of CBGs throughout the decade (Fig.
-S7b), implying a less mobile population.
-This is consistent with prior findings that residents of public housing and recipients of rental benefits have longer tenancies than average New York City residents [69, 71].
-To get a sense of the residents being served by NYCHA and their migration trajectories, we examine the demographics of the CBGs where NYCHA in-movers move from (Fig.
-S7c) and where NYCHA out-movers move to (Fig.
-S7d).
+NYCHA CBGs have had consistently lower in-migration and out-migration rates than the two comparison groups of CBGs throughout the decade (Supplementary Figure 2b), implying a less mobile population.
+This is consistent with prior findings that residents of public housing and recipients of rental benefits have longer tenancies than average New York City residents nycha_factsheet_2024, chen_measuring_2023.
+To get a sense of the residents being served by NYCHA and their migration trajectories, we examine the demographics of the CBGs where NYCHA in-movers move from (Supplementary Figure 2c) and where NYCHA out-movers move to (Supplementary Figure 2d).
 Compared to non-NYCHA in-movers, NYCHA in-movers are likelier to move from plurality Black or Hispanic CBGs, as well as bottom income quartile CBGs.
@@
 There is a significant amount of shuffling within the NYCHA system: 14% of in-movers to a NYCHA CBG came from another NYCHA CBG, and 13% of out-movers from a NYCHA building go to another NYCHA CBG.
-This shuffling may be due in part to mandated renovations or building closures that result in temporary resident relocations by NYCHA [72, 73, 74].
+This shuffling may be due in part to mandated renovations or building closures that result in temporary resident relocations by NYCHA navarro_harlem_2014, fernandez_new_2010, zaveri_improve_2023.
 On average across the decade that we study, NYCHA out-movers are likelier than non-NYCHA movers to remain within New York City (74% vs.
@@
 Overall, this analysis highlights how MIGRATE can provide insights into which residents are served by public housing policy.
-Beyond this case study of New York City, this approach could be extended to study public housing in other cities, or to study the effects of public housing foreclosures—which has been a topic of interest among other migration researchers [18].
-Appendix E Privacy Protections
+Beyond this case study of New York City, this approach could be extended to study public housing in other cities, or to study the effects of public housing foreclosures—which has been a topic of interest among other migration researchers phillips_measuring_2020.
+Appendix D Privacy Protections
 In this section we detail the measures we took to preserve the privacy of individuals in the MIGRATE data release.
@@
 This level of data aggregation is consistent with other data sources which have been made available to researchers.
-For example, mobility data often contains the aggregate counts of individuals moving from a given CBG to a given point of interest on a given day [75, 25, 76, 77, 78].
+For example, mobility data often contains the aggregate counts of individuals moving from a given CBG to a given point of interest on a given day abbiasov202415, chang2021mobility, kostandova2024systematic, xu2025using, yabe2024behaviour.
 Second, the matrices we release are probabilistic and rescaled: in particular, the fact that a particular entry of the CBG-CBG migration matrix is non-zero does not guarantee that a given number of individuals moved between those CBGs (or indeed, that any did at all).
 This is because individuals are mapped probabilistically to CBGs (e.g., if we have only a ZIP code for an individual, they are mapped probabilistically to CBGs within that ZIP code) and the migration matrices are rescaled multiple times.
-Third, we release the data only to researchers for non-profit use who agree to a data use agreement pledging not to re-identify individuals in the data, following manual review of their application and research project.
+Third, we release the data only to researchers for non-profit use who agree to a data use agreement pledging not to re-identify individuals in the data, and to adhere to privacy-protecting measures when storing data and presenting results, following manual review of their application and research project.
 Finally, we compute the average number of destinations (unique CBGs) for out-movers from each CBG.